### 20 Apr 2026 v2

# Rate-Distortion Theory for Deductive Sources under Closure Fidelity

Abstract
We study lossy compression of a finite statement source generated in a fixed deductive environment. The source symbols are statements in a knowledge base endowed with a proof system, and reconstruction fidelity is measured by preservation of deductive closure rather than by symbolwise equality. This induces, once the proof system and canonical order are fixed, a decomposition of the source into an irredundant core and redundant stored consequences. Under a natural disjointness condition on zero-distortion reconstruction sets, we show that the minimum zero-distortion rate equals the source mass of the core times the entropy of the source conditioned on that core. For reconstruction alphabets contained in the deductive closure of the source knowledge base, we further prove that the full rate–distortion function depends only on the core, so redundant states are invisible to both rate and distortion. When the decoder is limited to a bounded inference-depth budget, we obtain an exact rate–depth–distortion characterization. Under an additional order-robustness assumption identifying the chosen core with the order-free essential set, this characterization interpolates between classical symbolwise compression and unconstrained deductive compression. These results formulate deductive compression as a structured source-coding problem and quantify how shared inference structure changes the fundamental limits of communication.
I Introduction
Classical rate–distortion theory studies the minimum description rate required to represent a random source under a prescribed fidelity criterion [26, 7, 8]. In the standard formulation, the source alphabet is treated as a collection of symbols and distortion is imposed directly on source–reconstruction pairs. In many modern communication settings, however, the source is not naturally a flat alphabet but a finite knowledge base equipped with a shared proof system. In such settings, symbolwise reproduction can be unnecessarily stringent: if the receiver can reconstruct the same deductive consequences from a different stored representation, then the communication task has arguably been completed.
This observation suggests a source-coding problem in which fidelity is measured not by exact symbol reproduction but by preservation of deductive closure. We study this problem for a finite source set $S_{O}$ endowed with a fixed proof system and its induced closure operator $\text{Cn}(\cdot)$. The resulting source model carries an intrinsic redundancy structure: some states are irredundant premises, while others are stored consequences that can be re-derived once those premises are known. Once the proof system and canonical order are fixed, this induces a decomposition of the source into an irredundant core $A:=\text{Atom}(S_{O})$ and a redundant part $J:=S_{O}\setminus A$. Under closure fidelity, redundant states need not be transmitted faithfully, so the effective source rate can be strictly smaller than the classical rate.
Operationally, the source studied here is not a random knowledge base but a statement-level source generated within a fixed deductive environment. A finite knowledge base $S_{O}$, together with its proof system and closure operator, is fixed as part of the model, and the random source symbol is a statement $S_{o}\sim P_{O}$ drawn from that environment. Thus a source block models a stream of statement reports, updates, query answers, or semantic tokens produced relative to the same local knowledge base. This is why the distortion may depend on the ambient $S_{O}$ while remaining a standard single-letter distortion function.
The present formulation is related to, but distinct from, the broader semantic-communication and task-oriented communication literature, including early conceptual proposals[4], modern surveys[10], and recent mathematical formulations of semantic rate–distortion and semantic coding [18, 20, 32, 13, 19]. Our formulation is narrower and more Shannon-theoretic: we fix a finite deductive environment and let the proof system itself induce zero-distortion equivalence through closure preservation.
Related work and positioning.
Two classical information-theoretic lines are especially relevant. First, source coding with decoder side information and computation-oriented compression—from Slepian–Wolf and Wyner–Ziv to coding for computing—show how decoder-side structure or non-identity reconstruction goals can change fundamental compression limits [27, 31, 21]. Our setting differs in that the decoder’s advantage is not exogenous random side information and the reproduction target is not a fixed function specified in advance; instead, the shared proof system induces the fidelity criterion itself.
Second, the confusable-core phenomenon that appears here is closely connected to zero-error and graph-theoretic information theory. Shannon’s zero-error framework introduced confusability as an information-theoretic object, Körner’s graph entropy and Witsenhausen’s zero-error side-information formulation linked source coding to graph structure, and later work further clarified these connections [24, 16, 30, 2]. When Assumption 3.1 fails, our zero-distortion problem becomes a confusability problem on the irredundant core, which is why a graph-entropy-type completion is a natural next step.
A further adjacent line comes from logic and database theory, especially the fixpoint semantics of Horn programs, finite-domain Datalog, reasoning over materialized deductive stores, and the broader theory of incomplete information and certain answers [29, 5, 1, 9, 15, 12, 11]. Our redundant/materialized split is closest to the materialized-view perspective: the set $J$ behaves as stored consequences that need not be communicated once the irredundant premises are available. An incomplete-information analogue would replace deductive closure by a certain-answer operator, leading to a semantics-based fidelity notion. We do not pursue that extension here; the present paper remains a Shannon-theoretic study of a fixed deductive environment.
Finally, our work is also adjacent to recent rate–distortion formulations for semantic sources and to perception-constrained source coding [18, 32, 13, 23]. These works broaden the classical fidelity framework beyond symbolwise criteria, but they do not model a fixed deductive environment or exploit the proof-induced decomposition into an irredundant core and redundant stored consequences that is central to the present deductive-source model.
This paper asks three basic information-theoretic questions. First, what is the exact zero-distortion rate under closure fidelity? Second, how does the full rate–distortion function depend on the irredundant core? Third, how does this picture change when the decoder has only a limited inference-depth budget (a limited number of iterations)? We answer these questions in a finite-alphabet setting. Our sharpest converse statements require an explicit disjointness assumption on zero-distortion reconstruction sets; we state this assumption up front because a central goal of the paper is to identify precisely where exact characterization is possible and where additional structure is still needed.
Our main message is that closure fidelity converts deductive redundancy into rate savings. In particular, the zero-distortion problem reduces to coding only the irredundant core, and the full rate–distortion function decomposes into a contribution from that core alone. We further show that when the decoder can perform at most $\delta$ iterations of the immediate-consequence operator, the relevant effective source becomes a $\delta$-irredundant core. Under an additional order-robustness assumption aligning the chosen core with the order-free essential set, this yields an exact rate–depth–distortion interpolation between classical symbolwise compression and unconstrained deductive compression. Source–channel separation, closure-adapted converse bounds, and heterogeneous-receiver extensions then follow as consequences of this source-model characterization.
Main Contributions
  • A source-model formulation for deductive compression. We formulate a finite-alphabet rate–distortion problem in which the source is a knowledge base endowed with a fixed proof system and the fidelity criterion is preservation of deductive closure. This shifts the emphasis from scheme-specific semantic communication architectures to a source-model-level information-theoretic object.
  • An exact zero-distortion characterization under closure fidelity. Under Assumption 3.1, we prove that the minimum zero-distortion rate is
    $R_{\mathrm{sem}}(0)=P_{A}\,H(\pi_{A}),$
    where $A=\text{Atom}(S_{O})$ is the irredundant core, $P_{A}$ is its source mass, and $\pi_{A}$ is the source distribution conditioned on $A$ (Theorem 3.1). Thus the effective zero-distortion rate is determined by the core rather than by the full stored source.
  • A core-only decomposition of the full rate–distortion function. For reconstruction alphabets contained in $\text{Cn}(S_{O})$, we show that the entire rate–distortion function decomposes into a contribution from the irredundant core alone (Theorem 4.1). In this sense, redundant states are invisible to both rate and distortion.
  • An exact rate–depth–distortion law for limited inference. When the decoder is allowed only a bounded inference-depth budget, we introduce the $\delta$-irredundant core and prove the exact zero-distortion rate–depth law
    $R_{\mathrm{sem}}(0,\delta)=P_{\delta}\,H(\pi_{\delta}),$
    together with the corresponding full rate–depth–distortion function $R_{\mathrm{sem}}(D,\delta)$ over $(D,\delta)$ (Theorems 5.1 and 5.2). Under an additional order-robustness assumption ensuring $\text{Atom}(S_{O})=\operatorname{Ess}(S_{O})$, this yields a sharp interpolation between classical symbolwise coding and unconstrained deductive coding.
  • Information-theoretic consequences and extensions. The source characterization leads to source–channel separation statements, closure-adapted converse bounds of Fano type, and extensions to heterogeneous receivers with mismatched vocabularies. These results show how shared inference structure affects reliable communication beyond the homogeneous source-coding setting.
The remainder of the paper is organized as follows. Section II introduces the deductive source model and closure fidelity. Section III proves the exact zero-distortion rate. Section IV derives the full rate–distortion decomposition over the irredundant core. Section V studies bounded inference and the resulting rate–depth–distortion tradeoff. Section VI presents consequences including source–channel separation, converse bounds, and a heterogeneous extension. Appendix A collects deferred proofs of the main results in Sections III–VI. Appendix B records concrete Datalog classes satisfying the standing assumptions. Appendix C contains additional examples and counterexamples.
II Deductive Source Model and Closure Fidelity
This section isolates the minimum structure needed for the source-coding problem studied in this paper. At this stage no channel model is required. The basic object is a finite source alphabet of statements, equipped with a fixed proof system and its induced deductive closure operator. This structure induces, once the proof system and canonical order are fixed, a decomposition of the source into irredundant premises and redundant stored consequences, which in turn determines the closure-based fidelity criterion.
II-A Deductive Source and Irredundant Core
Assumption 2.1 (Fixed effective proof system).
We fix a finite ambient universe $\mathbb{S}_{O}$ of candidate statements together with an effective proof system $\mathsf{PS}$. For every finite $\Gamma\subseteq\mathbb{S}_{O}$ and every $s\in\mathbb{S}_{O}$, whether $\Gamma\vdash_{\mathrm{kb}}s$ is decidable. The induced deductive closure operator is
$\text{Cn}(\Gamma)\;:=\;\{s\in\mathbb{S}_{O}:\Gamma\vdash_{\mathrm{kb}}s\},$
viewed as the standard logical closure operator associated with the proof system [14]. It satisfies:
  1. (Cn1)
    Reflexivity: $\Gamma\subseteq\text{Cn}(\Gamma)$.
  2. (Cn2)
    Monotonicity: $\Gamma\subseteq\Gamma^{\prime}\Rightarrow\text{Cn}(\Gamma)\subseteq\text{Cn}(\Gamma^{\prime})$.
  3. (Cn3)
    Idempotence: $\text{Cn}(\text{Cn}(\Gamma))=\text{Cn}(\Gamma)$.
Assumption 2.2 (Finite source alphabet).
The transmitted knowledge base $S_{O}\subseteq\mathbb{S}_{O}$ is finite and effectively listable under a fixed canonical order. The reconstruction alphabet $\hat{S}_{O}\subseteq\mathbb{S}_{O}$ is also finite and nonempty.
Assumption 2.3 (Decidable redundancy test).
For every finite $\Gamma\subseteq S_{O}$ and every $s\in S_{O}$, membership $s\in\text{Cn}(\Gamma)$ is decidable.
Remark 2.1 (On Assumption 2.3).
Assumption 2.3 is implied by Assumption 2.1; it is stated separately only to highlight the specific decidability query used in the deletion procedure of Definition 2.2.
Definition 2.1 (Deductive source).
A deductive source is a pair $(S_{O},P_{O})$, where $S_{O}\subseteq\mathbb{S}_{O}$ is a finite knowledge base and $P_{O}\in\Delta(S_{O})$ is a probability distribution on $S_{O}$, together with the fixed proof system of Assumption 2.1.
Definition 2.2 (Irredundant core).
Let $S_{O}$ be a finite knowledge base. Define $\text{Atom}(S_{O})$ by the following deterministic deletion procedure: initialize $A\leftarrow S_{O}$; scan the elements of $S_{O}$ in the fixed canonical order; whenever the current element $s$ satisfies $s\in\text{Cn}(A\setminus\{s\})$, remove it from $A$. The final set is called the irredundant core of $S_{O}$.
Definition 2.3 (Core and redundant part).
For a deductive source $S_{O}$, write
$A:=\text{Atom}(S_{O}),\qquad J:=S_{O}\setminus A.$
The set $A$ is the irredundant core, and $J$ is the redundant part.
Proposition 2.1 (Core correctness).
Under Assumptions 2.12.3, the set $A:=\text{Atom}(S_{O})$ satisfies:
  1. (i)
    $\text{Cn}(A)=\text{Cn}(S_{O})$;
  2. (ii)
    for every $a\in A$, $a\notin\text{Cn}(A\setminus\{a\})$ (irredundancy);
  3. (iii)
    $\text{Atom}(S_{O})$ is uniquely determined by $S_{O}$, Cn, and the fixed canonical order.
Each deletion preserves closure. Indeed, if $s\in\text{Cn}(A\setminus\{s\})$, then $A\subseteq\text{Cn}(A\setminus\{s\})$, hence by monotonicity and idempotence,
$\text{Cn}(A)\subseteq\text{Cn}(\text{Cn}(A\setminus\{s\}))=\text{Cn}(A\setminus\{s\}),$
while the reverse inclusion is immediate from $A\setminus\{s\}\subseteq A$. By induction over the scan, the final output satisfies $\text{Cn}(A)=\text{Cn}(S_{O})$, proving (i).
For (ii), let $a\in A$ and let $A_{\mathrm{scan}}$ denote the current set when $a$ is scanned. Since $a$ is retained, $a\notin\text{Cn}(A_{\mathrm{scan}}\setminus\{a\})$. Later deletions only shrink the set, so $A\setminus\{a\}\subseteq A_{\mathrm{scan}}\setminus\{a\}$. Monotonicity then gives
$\text{Cn}(A\setminus\{a\})\subseteq\text{Cn}(A_{\mathrm{scan}}\setminus\{a\}),$
hence $a\notin\text{Cn}(A\setminus\{a\})$.
Part (iii) follows from the determinism of the deletion procedure once Cn and the canonical order are fixed. ∎
Remark 2.2 (Proof-system dependence and what is intrinsic).
The set $\text{Atom}(S_{O})$ is relative to the chosen proof system and, in general, to the canonical order. This dependence is not a defect of the model: it reflects the fact that communication value is measured relative to shared inference capability. What is primary in this paper is the induced rate–distortion quantity, not any one particular realizing core. Section III shows that the scalar zero-distortion rate is intrinsic to the deductive source model, even when different canonical orders may produce different irredundant cores. Appendix C-B then gives a sufficient condition under which the core itself becomes order-invariant.
II-B Closure Fidelity and Closure Distortion
The source model above suggests a fidelity criterion stronger than semantic similarity but weaker than symbolwise equality: a reconstruction should be considered faithful whenever it preserves the deductive consequences of the transmitted knowledge base.
Definition 2.4 (Closure fidelity).
For finite $S,\hat{S}\subseteq\mathbb{S}_{O}$, define
$\mathsf{F}_{\text{Cn}}(S,\hat{S})\;:=\;\frac{|\text{Cn}(S)\cap\text{Cn}(\hat{S})|}{|\text{Cn}(S)\cup\text{Cn}(\hat{S})|}$,
with $0/0:=1$. Thus $\mathsf{F}_{\text{Cn}}(S,\hat{S})=1$ if and only if $Cn(S^)\text$$\text{Cn}(\hat{S})$.
Remark 2.3 (Closure fidelity as a Jaccard-type similarity).
[…]
Remark 2.4 (Conceptual relation to semantic content).
[…]
Definition 2.5 (Hamming distortion).
The Hamming distortion is
$d_{H}(s_{o},\hat{s}_{o}):=\mathbf{1}[s_{o}\neq\hat{s}_{o}].$
Definition 2.6 (Closure distortion).
Fix a reference knowledge base $\Gamma\subseteq\mathbb{S}_{O}$ and write $\Gamma_{-s}:=\Gamma\setminus\{s\}$. For $s_{o}\in\Gamma$ and $\hat{s}_{o}\in\hat{S}_{O}$, define
$C_{s_{o}}:=\text{Cn}(\Gamma_{-s_{o}}\cup\{s_{o}\}),\qquad C_{\hat{s}_{o}}:=\text{Cn}(\Gamma_{-s_{o}}\cup\{\hat{s}_{o}\}),$
and set
$d_{\text{Cn}}(s_{o},\hat{s}_{o}\mid\Gamma):=1-\frac{|C_{s_{o}}\cap C_{\hat{s}_{o}}|}{|C_{s_{o}}\cup C_{\hat{s}_{o}}|},$ (2.1)
with $0/0:=0$. In the sequel we use the single-letter distortion $d_{\text{Cn}}(s_{o},\hat{s}_{o}):=d_{\text{Cn}}(s_{o},\hat{s}_{o}\mid S_{O})$.
Remark 2.5 (Why closure distortion is the right object here).
Under $d_{H}$, every symbol error is costly. Under $d_{\text{Cn}}$, replacing a statement by a different one may be free if the deductive closure is preserved. This is precisely the mechanism by which deductive redundancy becomes compressible. The entire zero-distortion theory developed in the next section rests on this distinction.
Proposition 2.2 (Redundant states are zero-distortion free).
Let $A=\text{Atom}(S_{O})$ and $J=S_{O}\setminus A$. If $j\in J$ and $\hat{s}\in\text{Cn}(S_{O})$, then
$d_{\text{Cn}}(j,\hat{s}\mid S_{O})=0.$
Because $A=\text{Atom}(S_{O})\subseteq S_{O}\setminus\{j\}$, one has by Proposition 2.1(i) and monotonicity that
$\text{Cn}(S_{O})=\text{Cn}(A)\subseteq\text{Cn}(S_{O}\setminus\{j\}).$
Since also $S_{O}\setminus\{j\}\subseteq S_{O}$, the reverse inclusion $\text{Cn}(S_{O}\setminus\{j\})\subseteq\text{Cn}(S_{O})$ is immediate. Hence
$\text{Cn}(S_{O}\setminus\{j\})=\text{Cn}(S_{O}).$
Because $\hat{s}\in\text{Cn}(S_{O})=\text{Cn}(S_{O}\setminus\{j\})$, one has
$(S_{O}\setminus\{j\})\cup\{\hat{s}\}\subseteq\text{Cn}(S_{O}\setminus\{j\}).$
By monotonicity and idempotence,
$\text{Cn}\bigl((S_{O}\setminus\{j\})\cup\{\hat{s}\}\bigr)=\text{Cn}(S_{O}\setminus\{j\})=\text{Cn}(S_{O}).$
Hence the two closures in Definition 2.6 coincide, and the Jaccard distance is zero. ∎
Definition 2.7 (Closure-based rate–distortion function).
For a deductive source $(S_{O},P_{O})$, a finite reconstruction alphabet $\hat{S}_{O}\subseteq\mathbb{S}_{O}$, and the single-letter distortion $d_{\text{Cn}}(\cdot,\cdot\mid S_{O})$, define
$R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid S}:\\ \mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})]\leq D\end{subarray}}I(\mathsf{S}_{o};\hat{\mathsf{S}}_{o}).$ (2.2)
This is the rate–distortion function of the deductive source under closure fidelity.
Remark 2.6 (Operational meaning of the source model).
The object being compressed in this paper is a statement stream drawn from a fixed deductive environment, not a random knowledge base. More precisely, a finite knowledge base $S_{O}$ is fixed once and for all as part of the source model, together with its proof system and closure operator, and the random source symbol $S_{o}$ is drawn from $P_{O}$ on $S_{O}$.
Operationally, one may think of $S_{o}$ as a statement report, update, query answer, or symbolic semantic token produced within the same local knowledge environment $S_{O}$. A block of length $m$ therefore models $m$ independent draws from the same deductive source environment. Accordingly, the distortion $d_{\text{Cn}}(s_{o},\hat{s}_{o}\mid S_{O})$ is a standard single-letter distortion function parameterized by the fixed ambient knowledge base $S_{O}$.
We do not study in this paper the distinct problem in which the knowledge base itself is random across blocks. That would lead to a higher-level hierarchical source model and is left for future work.
Remark 2.7 (Roadmap).
Section III studies the case $D=0$ and shows that the corresponding rate depends only on the irredundant core. Section V later replaces the unconstrained closure criterion by a bounded-inference version, yielding a rate–depth–distortion tradeoff.
III Exact Zero-Distortion Rate
This section gives the main zero-distortion characterization behind the paper. Throughout, let
$A:=\text{Atom}(S_{O}),\qquad J:=S_{O}\setminus A,\qquad P_{A}:=P_{O}(A),\qquad P_{J}:=1-P_{A},$
and, whenever $P_{A}>0$, let
$\pi_{A}(a):=\frac{P_{O}(a)}{P_{A}},\qquad a\in A.$
We adopt the convention $P_{A}H(\pi_{A}):=0$ when $P_{A}=0$.
The key technical issue is whether different core symbols can share the same zero-distortion reconstruction. When they cannot, the zero-distortion problem admits an exact single-letter solution.
Assumption 3.1 (Disjoint zero-distortion reconstruction sets).
For each $s\in S_{O}$, define its zero-distortion reconstruction set
$R_{\text{Cn}}(s):=\{\hat{s}_{o}\in\hat{S}_{O}:\ d_{\text{Cn}}(s,\hat{s}_{o}\mid S_{O})=0\}.$
We assume that for distinct $a_{1},a_{2}\in A$,
$R_{\text{Cn}}(a_{1})\cap R_{\text{Cn}}(a_{2})=\varnothing.$
Remark 3.1 (Fully redundant degenerate case).
If $A=\varnothing$ and there exists $\hat{s}^{\star}\in\hat{S}_{O}\cap\text{Cn}(S_{O})$, then the constant reconstruction $\hat{S}_{o}\equiv\hat{s}^{\star}$ achieves zero closure distortion for every source symbol, and hence
$R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})=0.$
The exact theorem below therefore focuses on the nondegenerate case $A\neq\varnothing$, where the core carries positive effective complexity.
Remark 3.2 (Zero-mass core case).
If $A\neq\varnothing$, $P_{A}=0$, and $A\subseteq\hat{S}_{O}$, then the source is supported on $J=S_{O}\setminus A$. Hence any fixed $a_{0}\in A$ used as a constant reconstruction achieves zero closure distortion almost surely, because $a_{0}\in A\subseteq\text{Cn}(S_{O})$ and every symbol in $J$ is zero-distortion free against every reconstruction in $\text{Cn}(S_{O})$ by Proposition 2.2. Therefore
$R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})=0.$
Accordingly, the only nontrivial case in Theorem 3.1 is $P_{A}>0$.
Remark 3.3 (Interpretation and naturality of Assumption 3.1).
Assumption 3.1 is a zero-distortion nonconfusability condition on the irredundant core. It is the closure-fidelity analogue of requiring pairwise disjoint zero-error decoding regions: a reconstruction that is closure-correct for one core symbol cannot simultaneously be closure-correct for a different core symbol.
Thus the assumption is not an arbitrary technical restriction but the precise condition under which the exact zero-distortion converse collapses to a single-letter entropy expression. When it fails, and still assuming $A\subseteq\hat{S}_{O}$, the achievability bound $R_{\mathrm{sem}}(0)\leq P_{A}H(\pi_{A})$ still holds, but the converse must account for overlap among the sets $\{R_{\text{Cn}}(a):a\in A\}$, leading to a genuine confusability problem on the core. Appendix B gives concrete Datalog classes in which the assumption is satisfied by construction.
Proposition 3.1 (Zero-distortion achievability).
Assume $P_{A}>0$ and $A\subseteq\hat{S}_{O}$. Then
$R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})\;\leq\;P_{A}\,H(\pi_{A}).$ (3.1)
Define the conditional distribution
$P_{\hat{S}\mid S}(\hat{s}_{o}\mid s_{o}):=\begin{cases}\mathbf{1}[\hat{s}_{o}=s_{o}],&s_{o}\in A,\\[3.0pt] \pi_{A}(\hat{s}_{o}),&s_{o}\in J,\end{cases}$
where $\pi_{A}$ is supported on $A\subseteq\hat{S}_{O}$.
For $a\in A$, the reconstruction is exact, so $d_{\text{Cn}}(a,a\mid S_{O})=0$. For $j\in J$, the reconstruction lies in $A\subseteq\text{Cn}(S_{O})$, hence $d_{\text{Cn}}(j,\hat{s}_{o}\mid S_{O})=0$ by Proposition 2.2. Therefore the expected distortion is zero.
Let $T$ be the auxiliary variable
$T(s_{o}):=\begin{cases}s_{o},&s_{o}\in A,\\ *,&s_{o}\in J.\end{cases}$
Under the above conditional law, $\hat{S}_{o}$ depends on $S_{o}$ only through $T$, and the output marginal is exactly $\pi_{A}$. Hence
$I(S_{o};\hat{S}_{o})=I(T;\hat{S}_{o})=H(\pi_{A})-P_{J}H(\pi_{A})=P_{A}\,H(\pi_{A}).$
This proves (3.1). ∎
Theorem 3.1 (Exact zero-distortion rate).
Assume $A\neq\varnothing$, $A\subseteq\hat{S}_{O}$, and Assumption 3.1. Then the closure-based zero-distortion rate is exactly
$R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})\;=\;P_{A}\,H(\pi_{A}),$ (3.2)
with the convention that the right-hand side is 0 when $P_{A}=0$.
If $P_{A}=0$, the claim follows from Remark 3.2. Assume henceforth that $P_{A}>0$.
The achievability part is Proposition 3.1. For the converse, collapse the source to the auxiliary variable
$T(s_{o}):=\begin{cases}s_{o},&s_{o}\in A,\\ *,&s_{o}\in J.\end{cases}$
Then $I(S_{o};\hat{S}_{o})\geq I(T;\hat{S}_{o})$ by data processing. Under zero closure distortion, each core symbol $a\in A$ must be reconstructed inside $R_{\text{Cn}}(a)$, and Assumption 3.1 makes these supports pairwise disjoint; symbols in $J$ impose no identifying constraint by Proposition 2.2. A disjoint-support entropy decomposition, combined with concavity of entropy, yields
$I(T;\hat{S}_{o})\geq P_{A}H(\pi_{A}).$
Together with Proposition 3.1, this proves (3.2). Full details are given in Appendix A-A. ∎
Remark 3.4 (Canonical-order dependence versus intrinsic rate).
The set $A=\text{Atom}(S_{O})$ may depend on the fixed canonical order used in Definition 2.2. This does not mean that the zero-distortion rate itself is order-dependent. Indeed, the left-hand side of (3.2), $R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})$, is defined entirely in terms of the source law $P_{O}$ and the closure distortion $d_{\text{Cn}}$, and therefore is intrinsic to the deductive source model. Consequently, for any canonical order under which Theorem 3.1 applies, the scalar $P_{A}H(\pi_{A})$ must equal the same intrinsic zero-distortion rate.
What may depend on the canonical order is the particular core/redundant decomposition that realizes this value. Appendix C-B gives a sufficient condition under which the core itself becomes order-invariant.
Corollary 3.1 (Uniform source and strict gain over Hamming fidelity).
Assume $A\neq\varnothing$, and let $k:=|A|$. Under the hypotheses of Theorem 3.1:
  1. (i)
    if $P_{O}$ is uniform on $S_{O}$, then
    $R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})=\frac{k}{|S_{O}|}\log k;$ (3.3)
  2. (ii)
    if, in addition, $S_{O}\subseteq\hat{S}_{O}$, $P_{O}$ has full support, and $J\neq\varnothing$, then
    $R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})<R(0;\,d_{H},\,P_{O})=H(P_{O}).$ (3.4)
For (i), under the uniform source, $P_{A}=k/|S_{O}|$ and $\pi_{A}$ is uniform on $A$, so $H(\pi_{A})=\log k$. Substituting into (3.2) gives (3.3).
For (ii), write $\pi_{J}$ for the source distribution conditioned on $J$. Then
$H(P_{O})=h_{b}(P_{A})+P_{A}H(\pi_{A})+P_{J}H(\pi_{J}).$
If $J\neq\varnothing$ and $P_{O}$ has full support, then $0<P_{A}<1$, so $h_{b}(P_{A})>0$. Hence
$H(P_{O})>P_{A}H(\pi_{A}).$
Using Theorem 3.1 and the classical identity $R(0;\,d_{H},\,P_{O})=H(P_{O})$ for zero Hamming distortion on a reconstruction alphabet containing $S_{O}$ [7, Sec. 10.5][8, Ch. 7], which is valid here because $S_{O}\subseteq\hat{S}_{O}$, completes the proof. ∎
Remark 3.5 (Interpretation).
Theorem 3.1 is the point at which deductive structure enters the information theory in a decisive way. Under Hamming fidelity, every source symbol must remain distinguishable and the zero-distortion rate is $H(P_{O})$. Under closure fidelity, only the irredundant core must remain distinguishable; the redundant part is absorbed into a single zero-cost component. This is why the exact answer is $P_{A}H(\pi_{A})$ rather than $H(P_{O})$.
Remark 3.6 (The main unresolved issue).
Assumption 3.1 identifies the main technical gap between the present theory and a fully general zero-distortion theory. Without disjointness, and still assuming $A\subseteq\hat{S}_{O}$, the upper bound
$R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})\leq P_{A}H(\pi_{A})$
still follows from Proposition 3.1, but the converse must quantify the overlap structure among the sets $R_{\text{Cn}}(a)$, $a\in A$. This leads naturally to a confusability graph on the core. A graph-entropy-type characterization of that general case is, in our view, the key next theoretical breakthrough needed to complete the zero-distortion picture. A related analysis for ambiguous alphabets under asymptotic equipartition is given in [28].
IV Core Decomposition of the Full Rate–Distortion Function
Section III identified the exact zero-distortion rate under closure fidelity. We now extend the analysis to arbitrary distortion levels. The main point is that the redundant part $J=S_{O}\setminus A$ remains invisible not only at $D=0$, but across the entire rate–distortion function. Thus the full problem decomposes into a classical rate–distortion problem on the irredundant core, scaled by the source mass of that core.
Definition 4.1 (Core sub-source rate–distortion function).
Let $A=\text{Atom}(S_{O})$ and let $\pi_{A}$ be the source distribution conditioned on $A$. Assume $\hat{S}_{O}\subseteq\text{Cn}(S_{O})\cap\mathbb{S}_{O}$. The core sub-source rate–distortion function is
$R^{(A)}(D^{\prime};\,d_{\text{Cn}},\,\pi_{A}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid A}:\\ \mathbb{E}[d_{\text{Cn}}(A_{o},\hat{S}_{o})]\leq D^{\prime}\end{subarray}}I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o}),$ (4.1)
where $\mathsf{A}_{o}\sim\pi_{A}$, the reconstruction alphabet is $\hat{S}_{O}$, and the distortion is $d_{\text{Cn}}(\cdot,\cdot\mid S_{O})$ restricted to inputs in $A$.
Theorem 4.1 (Core decomposition of the full rate–distortion function).
Assume $\hat{S}_{O}\subseteq\text{Cn}(S_{O})\cap\mathbb{S}_{O}$. Then, for every $D\geq 0$,
$R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})\;=\;P_{A}\cdot R^{(A)}\!\Bigl(\frac{D}{P_{A}};\,d_{\text{Cn}},\,\pi_{A}\Bigr),$ (4.2)
with the convention that the right-hand side is 0 when $P_{A}=0$.
In particular:
  1. (i)
    under the additional hypotheses $A\subseteq\hat{S}_{O}$ and Assumption 3.1, one has $R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})=P_{A}H(\pi_{A})$, recovering Theorem 3.1;
  2. (ii)
    $R_{\mathrm{sem}}(\cdot)$ is convex, non-increasing, and reducible to a finite-alphabet core problem that is amenable to standard numerical evaluation;
  3. (iii)
    redundant states are invisible to both rate and distortion.
The key observation is that every redundant state contributes zero closure distortion against every reconstruction in $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, so both the distortion constraint and the mutual-information minimization reduce to the irredundant core. A lower bound follows by introducing the auxiliary variable $T(s_{o})=s_{o}$ on $A$ and $T(s_{o})=*$ on $J$, and then applying entropy concavity exactly as in the converse of Theorem 3.1. Achievability is obtained by extending an optimal test channel for the core sub-source to the redundant part using a common output law. The full proof is given in Appendix A-B. ∎
Remark 4.1 (Disjointness is needed only for the exact zero-distortion formula).
The decomposition theorem itself does not require Assumption 3.1. That assumption is needed only to identify the core-only zero-distortion term in closed form as $H(\pi_{A})$. Thus the structural reduction to the core is more general than the exact zero-distortion evaluation.
Remark 4.2 (What remains nonclassical at positive distortion).
Theorem 4.1 reduces the full source to its core, but the reduced problem is still nonclassical because the distortion on the core is the closure distortion $d_{\text{Cn}}(\cdot,\cdot\mid S_{O})$, not a symbolwise distortion. Thus the theorem isolates the precise source of the semantic gain: redundancy disappears entirely, while the nontrivial positive-distortion geometry is concentrated on the core.
V Limited Inference and the Rate–Depth–Distortion Tradeoff
Sections III and IV treat the decoder as having unrestricted deductive power. We now introduce a bounded inference-depth budget. The resulting fidelity criterion distinguishes statements that are re-derivable eventually from those re-derivable within $\delta$ iterations of $T_{\mathsf{PS}}$. This yields a natural rate–depth–distortion tradeoff.
V-A Finite-Step Deduction and Derivation Depth
Assumption 5.1 (Finite-step closure dynamics).
For the proof systems considered in this section, there exists a computable monotone operator $T_{\mathsf{PS}}$ on finite subsets of $\mathbb{S}_{O}$ such that, for every finite $B\subseteq\mathbb{S}_{O}$,
  1. (T1)
    $T_{\mathsf{PS}}(B)\supseteq B$;
  2. (T2)
    $B\subseteq B^{\prime}\Rightarrow T_{\mathsf{PS}}(B)\subseteq T_{\mathsf{PS}}(B^{\prime})$;
  3. (T3)
    with $T_{\mathsf{PS}}^{0}(B):=B$ and $T_{\mathsf{PS}}^{n+1}(B):=T_{\mathsf{PS}}(T_{\mathsf{PS}}^{n}(B))$, one has
    $\text{Cn}(B)=\bigcup_{n\geq 0}T_{\mathsf{PS}}^{n}(B);$
  4. (T4)
    the sequence $T_{\mathsf{PS}}^{n}(B)$ stabilizes in finitely many steps.
This covers standard Datalog and Horn-clause settings over finite active domains [29, 5, 1, 9].
Definition 5.1 (Derivation depth).
For finite $B\subseteq\mathbb{S}_{O}$ and $s\in\mathbb{S}_{O}$, define
$\text{Dd}(s\mid B):=\min\{n\geq 0:\ s\in T_{\mathsf{PS}}^{n}(B)\},$
with $\text{Dd}(s\mid B):=\infty$ if $s\notin\text{Cn}(B)$.
Definition 5.2 (Maximum intrinsic derivation depth).
Let $A=\text{Atom}(S_{O})$. The maximum intrinsic derivation depth of $S_{O}$ is
$\mathsf{D_{d}}:=\max_{q\in S_{O}}\text{Dd}(q\mid A),$
with the convention $\max\varnothing:=0$.
Assumption 5.2 (Order-robust core for the bounded-inference section).
Throughout Section V, we additionally assume
$A=\text{Atom}(S_{O})=\operatorname{Ess}(S_{O}):=\{s\in S_{O}:\ s\notin\text{Cn}(S_{O}\setminus\{s\})\}.$
Equivalently, every element of the chosen core is globally nonderivable from the remaining stored source symbols. This assumption is satisfied, for example, under the essential-generation condition of Proposition C.1.
Remark 5.1 (Why the extra assumption is needed here).
The family $A_{\delta}$ is defined directly from derivability out of $S_{O}\setminus\{s\}$, so as $\delta$ increases it converges to the order-free essential set $\operatorname{Ess}(S_{O})$, not in general to an arbitrary order-dependent deletion core $\text{Atom}(S_{O})$. Assumption 5.2 is therefore exactly what ensures that the large-$\delta$ endpoint of the bounded-inference theory coincides with the zero-distortion law of Section III.
V-B Depth-Constrained Fidelity
Definition 5.3 ($\delta$-step closure distortion).
For integer $\delta\geq 0$, define
$d_{\text{Cn}}^{\delta}(s_{o},\hat{s}_{o}\mid S_{O}):=\begin{cases}0,&\text{if }S_{O}\subseteq T_{\mathsf{PS}}^{\delta}\bigl((S_{O}\setminus\{s_{o}\})\cup\{\hat{s}_{o}\}\bigr),\\[4.0pt] 1,&\text{otherwise}.\end{cases}$ (5.1)
Remark 5.2 (Interpretation of $d_{\text{Cn}}^{\delta}$).
The distortion $d_{\text{Cn}}^{\delta}$ is a bounded-inference fidelity criterion. It asks whether the entire source set $S_{O}$ can be reconstructed within at most $\delta$ iterations of the immediate-consequence operator after replacing $s_{o}$ by $\hat{s}_{o}$.
At $\delta=0$, one has
$d_{\text{Cn}}^{0}(s_{o},\hat{s}_{o}\mid S_{O})=\mathbf{1}[s_{o}\neq\hat{s}_{o}]=d_{H}(s_{o},\hat{s}_{o}).$
Thus the zero-inference endpoint recovers classical symbolwise fidelity.
Definition 5.4 ($\delta$-redundancy and $\delta$-irredundant core).
A state $s\in S_{O}$ is $\delta$-redundant if
$s\in T_{\mathsf{PS}}^{\delta}(S_{O}\setminus\{s\}).$
The corresponding $\delta$-irredundant core is
$A_{\delta}:=\text{Atom}_{\delta}(S_{O}):=\{s\in S_{O}:\ s\notin T_{\mathsf{PS}}^{\delta}(S_{O}\setminus\{s\})\}.$
Write
$P_{\delta}:=P_{O}(A_{\delta}),\qquad\pi_{\delta}(s):=\frac{P_{O}(s)}{P_{\delta}},\quad s\in A_{\delta},$
whenever $P_{\delta}>0$.
Lemma 5.1 ($\delta$-redundant states are distortion-free).
If $s\notin A_{\delta}$, then
$d_{\text{Cn}}^{\delta}(s,\hat{s}\mid S_{O})=0,\qquad\forall\,\hat{s}\in\hat{S}_{O}.$
If $s\notin A_{\delta}$, then by definition
$s\in T_{\mathsf{PS}}^{\delta}(S_{O}\setminus\{s\}).$
By monotonicity of $T_{\mathsf{PS}}^{\delta}$,
[…]
Hence both $S_{O}\setminus\{s\}$ and $s$ belong to the latter set, so the whole source set $S_{O}$ is recovered within $\delta$ steps. Therefore $d_{\text{Cn}}^{\delta}(s,\hat{s}\mid S_{O})=0$. ∎
Proposition 5.1 ($\delta$-core filtration).
The family $\{A_{\delta}\}_{\delta\geq 0}$ satisfies:
  1. (i)
    $A_{0}=S_{O}$;
  2. (ii)
    $A_{\delta+1}\subseteq A_{\delta}$ for every $\delta\geq 0$;
  3. (iii)
    under Assumption 5.2, $A_{\delta}=A$ for every $\delta\geq\mathsf{D_{d}}$.
Part (i) follows from $T_{\mathsf{PS}}^{0}(S_{O}\setminus\{s\})=S_{O}\setminus\{s\}$, which never contains $s$.
For (ii), monotonicity in the number of iterations gives
$T_{\mathsf{PS}}^{\delta}(S_{O}\setminus\{s\})\subseteq T_{\mathsf{PS}}^{\delta+1}(S_{O}\setminus\{s\}),$
so once $s$ becomes $\delta$-redundant, it remains $(\delta+1)$-redundant.
For (iii), let $j\in J=S_{O}\setminus A$. Since $\text{Cn}(A)=\text{Cn}(S_{O})$, one has $j\in\text{Cn}(A)$, and by the definition of $\mathsf{D_{d}}$,
$j\in T_{\mathsf{PS}}^{\mathsf{D_{d}}}(A).$
Because $A\subseteq S_{O}\setminus\{j\}$, monotonicity yields
$j\in T_{\mathsf{PS}}^{\mathsf{D_{d}}}(S_{O}\setminus\{j\}),$
hence $j\notin A_{\delta}$ for every $\delta\geq\mathsf{D_{d}}$.
Conversely, if $a\in A$, then by Assumption 5.2,
$a\notin\text{Cn}(S_{O}\setminus\{a\}).$
Therefore $a\notin T_{\mathsf{PS}}^{\delta}(S_{O}\setminus\{a\})$ for every finite $\delta$, because $T_{\mathsf{PS}}^{\delta}(S_{O}\setminus\{a\})\subseteq\text{Cn}(S_{O}\setminus\{a\})$. Hence $a\in A_{\delta}$ for all $\delta$, proving $A_{\delta}=A$ when $\delta\geq\mathsf{D_{d}}$. ∎
Assumption 5.3 (Disjoint zero-distortion sets for the $\delta$-core).
For each $\delta\geq 0$ and each $a\in A_{\delta}$, define
$R_{\delta}(a):=\{\hat{s}_{o}\in\hat{S}_{O}:\ d_{\text{Cn}}^{\delta}(a,\hat{s}_{o}\mid S_{O})=0\}.$
We assume that for distinct $a_{1},a_{2}\in A_{\delta}$,
$R_{\delta}(a_{1})\cap R_{\delta}(a_{2})=\varnothing.$
V-C Exact Rate–Depth–Distortion Laws
Definition 5.5 (Rate–depth–distortion function).
For $\delta\geq 0$, define
$R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid S}:\\ \mathbb{E}[d_{\text{Cn}}^{\delta}(S_{o},\hat{S}_{o})]\leq D\end{subarray}}I(\mathsf{S}_{o};\hat{\mathsf{S}}_{o}).$ (5.2)
Theorem 5.1 (Exact zero-distortion rate–depth tradeoff).
Fix $\delta\geq 0$. Assume $A_{\delta}\subseteq\hat{S}_{O}$ and that Assumption 5.3 holds at this $\delta$. Then
$R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})\;=\;P_{\delta}\,H(\pi_{\delta}),$ (5.3)
with the convention that the right-hand side is 0 when $P_{\delta}=0$.
The proof is the exact analogue of Theorem 3.1, with $A$, $P_{A}$, and $\pi_{A}$ replaced by $A_{\delta}$, $P_{\delta}$, and $\pi_{\delta}$, and with Lemma 5.1 playing the role of Proposition 2.2.
Indeed, define
$T_{\delta}(s_{o}):=\begin{cases}s_{o},&s_{o}\in A_{\delta},\\ *,&s_{o}\notin A_{\delta}.\end{cases}$
Under zero $\delta$-distortion, every $a\in A_{\delta}$ must be reconstructed inside its zero-distortion set $R_{\delta}(a)$, and these sets are pairwise disjoint by Assumption 5.3. For $s_{o}\notin A_{\delta}$, the output is unconstrained by Lemma 5.1. The same disjoint-support entropy identity as in the proof of Theorem 3.1 therefore yields the converse
$I(S_{o};\hat{S}_{o})\geq P_{\delta}H(\pi_{\delta}).$
Achievability is obtained by reproducing each $a\in A_{\delta}$ exactly and mapping every $s_{o}\notin A_{\delta}$ to a common distribution $\pi_{\delta}$ supported on $A_{\delta}$. Then zero $\delta$-distortion holds, and the resulting mutual information is exactly $P_{\delta}H(\pi_{\delta})$. ∎
Corollary 5.1 (Endpoints and monotonicity of the zero-distortion rate–depth function).
Assume $S_{O}\subseteq\hat{S}_{O}$, that Assumption 5.3 holds for every $\delta\geq 0$, and that Assumption 5.2 holds. Then:
  1. (i)
    $R_{\mathrm{sem}}(0,0;\,d_{\text{Cn}}^{0},\,P_{O})=H(P_{O})$;
  2. (ii)
    $R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})=P_{A}H(\pi_{A})$ for all $\delta\geq\mathsf{D_{d}}$;
  3. (iii)
    the function $\delta\mapsto R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})$ is non-increasing.
Part (i) is Theorem 5.1 at $\delta=0$, where $A_{0}=S_{O}$, $P_{0}=1$, and $\pi_{0}=P_{O}$. Part (ii) follows from Proposition 5.1(iii) and Theorem 5.1. Indeed, for every $\delta\geq\mathsf{D_{d}}$, one has $A_{\delta}=A$, hence $P_{\delta}=P_{A}$ and $\pi_{\delta}=\pi_{A}$. Since $S_{O}\subseteq\hat{S}_{O}$, also $A_{\delta}=A\subseteq\hat{S}_{O}$, so the hypotheses of Theorem 5.1 are satisfied at each such $\delta$. Therefore
$R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})=P_{A}H(\pi_{A}),\qquad\delta\geq\mathsf{D_{d}}.$
Part (iii) holds because the zero $\delta$-distortion feasible set enlarges with $\delta$, so the minimum mutual information cannot increase. ∎
Definition 5.6 ($\delta$-core sub-source rate–distortion function).
For $\delta\geq 0$ and $P_{\delta}>0$, define
$R^{(A_{\delta})}(D^{\prime};\,d_{\text{Cn}}^{\delta},\,\pi_{\delta}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid A_{\delta}}:\\ \mathbb{E}[d_{\text{Cn}}^{\delta}(A_{\delta,o},\hat{S}_{o})]\leq D^{\prime}\end{subarray}}I(\mathsf{A}_{\delta,o};\hat{\mathsf{S}}_{o}),$ (5.4)
where $\mathsf{A}_{\delta,o}\sim\pi_{\delta}$, the reconstruction alphabet is $\hat{S}_{O}$, and the distortion is $d_{\text{Cn}}^{\delta}(\cdot,\cdot\mid S_{O})$ restricted to inputs in $A_{\delta}$.
Theorem 5.2 (Full rate–depth–distortion decomposition).
For every $\delta\geq 0$,
$R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})\;=\;P_{\delta}\cdot R^{(A_{\delta})}\!\Bigl(\frac{D}{P_{\delta}};\,d_{\text{Cn}}^{\delta},\,\pi_{\delta}\Bigr),$ (5.5)
with the convention that the right-hand side is 0 when $P_{\delta}=0$, where $R^{(A_{\delta})}$ is the rate–distortion function of the $\delta$-core sub-source $(A_{\delta},\pi_{\delta})$ under the distortion $d_{\text{Cn}}^{\delta}(\cdot,\cdot\mid S_{O})$ restricted to inputs in $A_{\delta}$.
In particular, at $\delta=0$, since $A_{0}=S_{O}$ and $d_{\text{Cn}}^{0}=d_{H}$, one recovers the ordinary Hamming rate–distortion problem on source alphabet $S_{O}$ and reconstruction alphabet $\hat{S}_{O}$:
$R_{\mathrm{sem}}(D,0;\,d_{\text{Cn}}^{0},\,P_{O})=\min_{\begin{subarray}{c}P_{\hat{S}\mid S}:\\ \mathbb{E}[d_{H}(S_{o},\hat{S}_{o})]\leq D\end{subarray}}I(\mathsf{S}_{o};\hat{\mathsf{S}}_{o}).$
By Lemma 5.1, every symbol outside $A_{\delta}$ is distortion-free under $d_{\text{Cn}}^{\delta}$, so both the distortion constraint and the mutual-information minimization reduce to the $\delta$-core. The lower bound follows by introducing $T_{\delta}(s_{o})=s_{o}$ on $A_{\delta}$ and $T_{\delta}(s_{o})=*$ on its complement, exactly as in the proof of Theorem 4.1. Achievability is obtained by extending an optimal test channel for the $\delta$-core sub-source to the $\delta$-redundant part using a common output law. At $\delta=0$, one has $A_{0}=S_{O}$ and $d_{\text{Cn}}^{0}=d_{H}$, so the formula reduces to the classical Hamming rate–distortion function [25, 7, 8]. Full details are given in Appendix A-E. ∎
Remark 5.3 (Interpretation).
The family $\{A_{\delta}\}_{\delta\geq 0}$ provides a quantitative depth-for-rate exchange. As $\delta$ increases, more states become reconstructible within the allowed depth budget and therefore disappear from the effective source. At $\delta=0$, nothing is free and the theory collapses to classical source coding. At $\delta\geq\mathsf{D_{d}}$, all eventually derivable shortcuts become free and, under the standing assumptions guaranteeing the large-$\delta$ endpoint identification, one recovers the deductive compression law of Section III.
Remark 5.4 (Endpoint caution).
The bounded-inference distortion $d_{\text{Cn}}^{\delta}$ is a binary feasibility distortion, whereas the unconstrained closure distortion $d_{\text{Cn}}$ of Sections II and IV is a Jaccard-type distortion. Thus Theorem 5.2 should be interpreted as the exact bounded-inference analogue of the rate–distortion function. Its zero-distortion endpoint agrees with Theorem 3.1, but for $D>0$ it is a distinct operational distortion family.
VI Consequences: Separation, Converse Bounds, and One Heterogeneous Extension
The preceding sections are source-coding results. We now record three consequences. First, because the relevant distortions are single-letter and bounded, the classical source–channel separation theorem applies verbatim [25, 7, 8, 22]. Second, the core/redundant decomposition yields sharper converse bounds than symbolwise fidelity. Third, the same zero-distortion mechanism extends to a restricted reconstruction alphabet, giving a compact heterogeneous-receiver extension.
Remark 6.1 (Degenerate core cases).
If $A=\varnothing$ and $\hat{S}_{O}\cap\text{Cn}(S_{O})\neq\varnothing$, then a constant reconstruction to any $\hat{s}^{\star}\in\hat{S}_{O}\cap\text{Cn}(S_{O})$ achieves zero closure distortion, so the effective zero-distortion rate is 0. If $|A|=1$, then all entropy and blocklength expressions involving the core reduce to their obvious one-symbol values, and any term of the form $\epsilon\log(|A|-1)$ should be read through the trivial one-symbol case rather than literally as a logarithm. Accordingly, the nontrivial converse and operational bounds below are primarily of interest when $|A|\geq 2$.
VI-A Separation and Operational Blocklength
Remark 6.2 (Two operational regimes).
The results in this section use two standard operational viewpoints. In the source-coding regime, one transmits an i.i.d. block of statement symbols drawn from the fixed deductive source $(S_{O},P_{O})$, leading to rate–distortion and separation statements. In the message-set regime, one treats elements of a finite set $\mathcal{M}\subseteq S_{O}$ as equiprobable messages and studies the minimum blocklength needed for reliable transmission. The former is the asymptotic rate–distortion viewpoint; the latter is the finite-message operational viewpoint underlying the converse and achievability bounds below.
Definition 6.1 (Block code and reliability criteria).
Let $W:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be a discrete memoryless channel with Shannon capacity $C(W)$. An $(n,M)$ block code for message set $\mathcal{M}\subseteq S_{O}$ consists of an encoder $f_{n}:\mathcal{M}\to\mathcal{X}^{n}$ and a decoder $g_{n}:\mathcal{Y}^{n}\to\hat{S}_{O}$, where $|\mathcal{M}|=M$.
When message $m\in\mathcal{M}$ is sent, write $\hat{\mathsf{S}}_{o}^{(m)}:=g_{n}(\hat{\mathsf{Y}}^{n})$ for the decoder output. The Hamming error probability and closure error probability are
$P_{e}^{(n)}$ $:=\max_{m\in\mathcal{M}}\Pr[\hat{\mathsf{S}}_{o}^{(m)}\neq m],$ (6.1)
$P_{e,\text{Cn}}^{(n)}$ $:=\max_{m\in\mathcal{M}}\Pr[d_{\text{Cn}}(m,\hat{\mathsf{S}}_{o}^{(m)}\mid S_{O})>0].$ (6.2)
Remark 6.3 (Block source interpretation).
In the separation results below, a length-$m$ source block means $m$ i.i.d. draws from the same deductive source $(S_{O},P_{O})$, with the knowledge base $S_{O}$ fixed throughout the block. Thus the channel-coding consequences are standard source–channel consequences for a structured but stationary finite-alphabet source.
Theorem 6.1 (Deductive source–channel separation).
Let $W$ be a discrete memoryless channel with capacity $C(W)$. Let an i.i.d. block of $m$ source symbols from $(S_{O},P_{O})$ be transmitted over $n$ channel uses, and write $\kappa:=n/m$.
For every bounded single-letter distortion family considered in this paper, the classical separation principle applies:
  1. (i)
    if
    $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})<\kappa\,C(W),$
    then distortion level $D$ under closure distortion is achievable;
  2. (ii)
    if a distortion level $D$ under closure distortion is achievable, then
    $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})\leq\kappa\,C(W);$
  3. (iii)
    for every fixed $\delta\geq 0$, if
    $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})<\kappa\,C(W),$
    then distortion level $D$ under the bounded-inference distortion $d_{\text{Cn}}^{\delta}$ is achievable; conversely, if that distortion level is achievable, then
    $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})\leq\kappa\,C(W).$
In particular, in the nondegenerate case $A\neq\varnothing$, whenever the hypotheses of Theorem 3.1 for $d_{\text{Cn}}$ and of Theorem 5.1 for the relevant $\delta$ are satisfied, one has:
$P_{A}H(\pi_{A})<\kappa\,C(W)$ $\;\Longrightarrow\;\text{zero closure distortion is achievable,}$ (6.3)
$\text{if zero closure distortion is achievable, then}\qquad P_{A}H(\pi_{A})$ $\leq\kappa\,C(W),$ (6.4)
$P_{\delta}H(\pi_{\delta})<\kappa\,C(W)$ $\;\Longrightarrow\;\text{zero $\delta$-distortion is achievable,}$ (6.5)
$\text{if zero $\delta$-distortion is achievable, then}\qquad P_{\delta}H(\pi_{\delta})$ $\leq\kappa\,C(W).$ (6.6)
Each distortion family considered here is a bounded single-letter distortion on a finite alphabet. Therefore the standard source–channel separation theorem for discrete memoryless sources and channels applies; see, e.g., [7, Chs. 7 and 10], [8, Ch. 3], [22, Chs. 7 and 8]. For $d_{\text{Cn}}$, the relevant source-coding function is $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$; for $d_{\text{Cn}}^{\delta}$, it is $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})$. The zero-distortion thresholds follow by substituting the exact expressions from Theorems 3.1 and 5.1. Full details are given in Appendix A-F. ∎
Corollary 6.1 (Depth-budget thresholds from separation).
Assume $S_{O}\subseteq\hat{S}_{O}$, Assumption 5.2, and Assumption 5.3 for every $\delta\geq 0$. Fix $\kappa=n/m$ and a channel $W$ with capacity $C(W)$. Define
$\delta_{\mathrm{ach}}^{*}(S_{O},P_{O},W,\kappa)$ $:=\min\{\delta\geq 0:\ P_{\delta}H(\pi_{\delta})<\kappa C(W)\},$ (6.7)
$\delta_{\mathrm{nec}}^{*}(S_{O},P_{O},W,\kappa)$ $:=\min\{\delta\geq 0:\ P_{\delta}H(\pi_{\delta})\leq\kappa C(W)\},$ (6.8)
with the convention $\min\varnothing:=\infty$.
Then:
  1. (i)
    every $\delta\geq\delta_{\mathrm{ach}}^{*}$ is sufficient for zero $\delta$-distortion transmission;
  2. (ii)
    if zero $\delta$-distortion transmission is achievable, then necessarily $\delta\geq\delta_{\mathrm{nec}}^{*}$;
  3. (iii)
    one always has
    $\delta_{\mathrm{nec}}^{*}\leq\delta_{\mathrm{ach}}^{*};$
  4. (iv)
    if no $\delta$ satisfies $P_{\delta}H(\pi_{\delta})=\kappa C(W)$, then
    $\delta_{\mathrm{ach}}^{*}=\delta_{\mathrm{nec}}^{*};$
  5. (v)
    if $H(P_{O})<\kappa C(W)$, then
    $\delta_{\mathrm{ach}}^{*}=\delta_{\mathrm{nec}}^{*}=0;$
  6. (vi)
    if $P_{A}H(\pi_{A})>\kappa C(W)$, then no finite depth budget suffices;
  7. (vii)
    if $P_{A}H(\pi_{A})<\kappa C(W)<H(P_{O})$, then
    $1\leq\delta_{\mathrm{nec}}^{*}\leq\delta_{\mathrm{ach}}^{*}\leq\mathsf{D_{d}}.$
By Corollary 5.1, the function
$\phi(\delta):=P_{\delta}H(\pi_{\delta})$
is non-increasing in $\delta$ and stabilizes at $P_{A}H(\pi_{A})$ for all $\delta\geq\mathsf{D_{d}}$. Part (i) follows from Theorem 6.1(iii) together with the definition of $\delta_{\mathrm{ach}}^{*}$. Part (ii) follows from the converse direction in Theorem 6.1(iii) and the definition of $\delta_{\mathrm{nec}}^{*}$. Part (iii) is immediate because
$\{\delta:\phi(\delta)<\kappa C(W)\}\subseteq\{\delta:\phi(\delta)\leq\kappa C(W)\}.$
Part (iv) follows because when equality never occurs, the two sets above coincide. Part (v) follows from $\phi(0)=H(P_{O})$. Part (vi) follows because $\phi(\delta)=P_{A}H(\pi_{A})$ for all $\delta\geq\mathsf{D_{d}}$, so if $P_{A}H(\pi_{A})>\kappa C(W)$, even the stabilized large-$\delta$ endpoint lies above channel capability. Part (vii) follows because $\phi(0)=H(P_{O})>\kappa C(W)$, so neither threshold is attained at $\delta=0$, whereas $\phi(\mathsf{D_{d}})=P_{A}H(\pi_{A})<\kappa C(W)$; monotonicity of $\phi$ then gives
$1\leq\delta_{\mathrm{nec}}^{*}\leq\delta_{\mathrm{ach}}^{*}\leq\mathsf{D_{d}}.$
Theorem 6.2 (Message-set converse bounds).
Let $A=\text{Atom}(S_{O})$. For any $(n,M)$ block code over a discrete memoryless channel $W$ with capacity $C(W)$:
  1. (i)
    if $P_{e}^{(n)}\leq\epsilon$, then
    $\log M\leq\frac{nC(W)+1}{1-\epsilon};$ (6.9)
  2. (ii)
    if $A\neq\varnothing$, $\mathcal{M}=S_{O}$, $P_{e,\text{Cn}}^{(n)}\leq\epsilon$, and Assumption 3.1 holds, then
    $\log|A|\leq\frac{nC(W)+1}{1-\epsilon}.$ (6.10)
Part (i) is the standard Fano converse [7, Sec. 2.10][22, Sec. 1.9].
For (ii), let $M_{A}$ be uniformly distributed on $A$ and consider the induced $|A|$-message subproblem. Since $P_{e,\text{Cn}}^{(n)}\leq\epsilon$ in the max-error sense, the average closure error probability for $M_{A}$ is also at most $\epsilon$. Under Assumption 3.1, the zero-distortion reconstruction sets $\{R_{\text{Cn}}(a):a\in A\}$ are pairwise disjoint, so a decoder output that is closure-correct for one core symbol cannot be closure-correct for another. Thus closure-correct decoding of core symbols induces an ordinary $|A|$-ary message recovery problem, and Fano’s inequality yields (6.10). ∎
Theorem 6.3 (Closure-reliable achievability).
Let $W$ be a discrete memoryless channel with capacity $C(W)$. Then:
  1. (i)
    if $S_{O}\subseteq\hat{S}_{O}$, then the full source $S_{O}$ can be transmitted Hamming-reliably whenever
    $\frac{\log|S_{O}|}{n}<C(W);$
  2. (ii)
    if $A\neq\varnothing$ and $A\subseteq\hat{S}_{O}$, then the full source $S_{O}$ can be transmitted closure-reliably whenever
    $\frac{\log|A|}{n}<C(W).$
Part (i) is Shannon channel coding for a message set of size $|S_{O}|$; see, e.g., [26, 7, 8].
For (ii), code only the core $A$. Choose an $(n,|A|)$ reliable channel code for $A$. If the transmitted symbol is $a\in A$, decode it as $a$. If the transmitted symbol is $j\in J$, encode it using the codeword of any fixed $a_{0}\in A$, and decode to whatever core symbol the channel decoder outputs. Closure correctness for $a\in A$ follows from correct decoding, and closure correctness for $j\in J$ follows from Proposition 2.2. ∎
Remark 6.4 (Heuristic operational deductive compression ratio).
Assume $A\neq\varnothing$, Assumption 3.1, and $S_{O}\subseteq\hat{S}_{O}$. Ignoring finite-blocklength backoff terms and the additive constant in Fano’s inequality, Theorems 6.2 and 6.3 suggest the benchmark blocklengths
$n_{H}^{\dagger}:=\left\lceil\frac{\log|S_{O}|}{C(W)}\right\rceil,\qquad n_{\text{Cn}}^{\dagger}:=\left\lceil\frac{\log|A|}{C(W)}\right\rceil.$
Accordingly, the corresponding heuristic deductive compression ratio is
$\frac{n_{\text{Cn}}^{\dagger}}{n_{H}^{\dagger}}\approx\frac{\log|A|}{\log|S_{O}|},$
which is strictly smaller than $1$ whenever $J\neq\varnothing$. This interpretation is asymptotic and operational, rather than an exact finite-blocklength theorem.
VI-B Closure-Adapted Converse Bounds
Theorem 6.4 (Closure-adapted Fano-type lower bound).
Let $(S_{O},P_{O})$ be full-support, and assume that the reconstruction alphabet satisfies
$\hat{S}_{O}\subseteq\text{Cn}(S_{O})\cap\mathbb{S}_{O}.$
Let $\hat{\mathsf{S}}_{o}$ be any reconstruction random variable on $\hat{S}_{O}$, and define
$\epsilon_{\text{Cn}}:=\Pr[d_{\text{Cn}}(S_{o},\hat{S}_{o}\mid S_{O})>0].$
If $A=\varnothing$, then $\epsilon_{\text{Cn}}=0$ and the conclusion is trivial. Otherwise,
$\epsilon_{\text{Cn}}=\sum_{a\in A}P_{O}(a)\,\Pr[\hat{S}_{o}\notin R_{\text{Cn}}(a)\mid S_{o}=a].$
Under Assumption 3.1,
$I(S_{o};\hat{S}_{o})\;\geq\;P_{A}\,H(\pi_{A})-h_{b}(\epsilon_{\text{Cn}})-\epsilon_{\text{Cn}}\log(|A|-1),$ (6.11)
for $|A|\geq 2$, while for $|A|=1$ the bound reduces to $I(S_{o};\hat{S}_{o})\geq 0$.
When $J=\varnothing$, the inequality takes the same Fano form on the full source alphabet $S_{O}$; if, in addition, $R_{\text{Cn}}(s)=\{s\}$ for every $s\in S_{O}$, it is exactly the classical Fano inequality.
Because $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, every redundant input is automatically closure-correct by Proposition 2.2, so closure errors can arise only on the core $A$. Conditioning on the event $B:=\mathbf{1}[S_{o}\in A]$, one reduces the problem to a $|A|$-ary source with law $\pi_{A}$. Under Assumption 3.1, the sets $\{R_{\text{Cn}}(a):a\in A\}$ are pairwise disjoint, so the decoder output induces an ordinary decision rule on $A\cup\{\mathsf{e}\}$. Applying the classical Fano inequality to the conditioned core source and then multiplying by $P_{A}$ yields the bound. Full details are given in Appendix A-G. ∎
Remark 6.5 (Why the bound is sharper than the classical one).
The reference term in (6.11) is $P_{A}H(\pi_{A})$, not $H(P_{O})$. Thus the converse is aligned with the effective source complexity under closure fidelity rather than with the raw symbol entropy of the stored source. This is precisely the right notion of converse once redundant states are free.
VI-C One Heterogeneous Extension
We now consider a receiver whose reconstruction alphabet is a restricted subset $V\subseteq\mathbb{S}_{O}$. This captures vocabulary mismatch or heterogeneous receivers. The question is whether the exact zero-distortion rate remains the same when the decoder is no longer allowed to reproduce arbitrary elements of the sender’s alphabet.
Definition 6.2 (Restricted-alphabet semantic rate–distortion function).
For $V\subseteq\mathbb{S}_{O}$, define
$R_{\mathrm{sem}}^{(V)}(D;\,d_{\text{Cn}},\,P_{O}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid S}:\\ \hat{S}_{o}\in V\ \text{a.s.}\\ \mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})]\leq D\end{subarray}}I(\mathsf{S}_{o};\hat{\mathsf{S}}_{o}).$ (6.12)
Remark 6.6 (Fully redundant restricted-alphabet case).
If $A=\varnothing$ and $V\cap\text{Cn}(S_{O})\neq\varnothing$, then a constant reconstruction to any $\hat{s}^{\star}\in V\cap\text{Cn}(S_{O})$ achieves zero closure distortion, so
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=0.$
The theorem below therefore focuses on the nondegenerate case $A\neq\varnothing$.
Theorem 6.5 (Heterogeneous zero-distortion extension).
Let $(S_{O},P_{O})$ be a deductive source with nonempty core $A=\text{Atom}(S_{O})$, and let $V\subseteq\mathbb{S}_{O}$ be a restricted reconstruction alphabet. For each $a\in A$, define
$R_{\text{Cn}}^{(V)}(a):=\{\hat{s}\in V:\ d_{\text{Cn}}(a,\hat{s}\mid S_{O})=0\}.$
Assume that:
  1. (H1)
    $R_{\text{Cn}}^{(V)}(a)\neq\varnothing$ for every $a\in A$;
  2. (H2)
    for distinct $a_{1},a_{2}\in A$,
    $R_{\text{Cn}}^{(V)}(a_{1})\cap R_{\text{Cn}}^{(V)}(a_{2})=\varnothing.$
Then
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=P_{A}\,H(\pi_{A}),$ (6.13)
with the convention that the right-hand side is 0 when $P_{A}=0$.
Conversely, if there exists $a\in A$ with $P_{O}(a)>0$ and $R_{\text{Cn}}^{(V)}(a)=\varnothing$, then the zero-distortion feasible set is empty, i.e.,
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=\infty.$
If $P_{A}=0$, the claim is immediate by constant reconstruction to any representative of a core symbol guaranteed by (H1). Assume henceforth that $P_{A}>0$.
Choose, for each $a\in A$, one representative $\phi(a)\in R_{\text{Cn}}^{(V)}(a)$. Assumption (H2) makes $\phi$ injective, so the core can be coded exactly as in Proposition 3.1, while the redundant part can be mapped to a common distribution supported on $\phi(A)$. This gives zero distortion and mutual information $P_{A}H(\pi_{A})$.
For the converse, any zero-distortion test channel supported on $V$ must map each $a\in A$ into $R_{\text{Cn}}^{(V)}(a)$, and these sets are pairwise disjoint by (H2). Thus the converse proof of Theorem 3.1 applies verbatim after replacing $R_{\text{Cn}}(a)$ by $R_{\text{Cn}}^{(V)}(a)$. Full details are given in Appendix A-H. ∎
Corollary 6.2 (Core coverage as a simple sufficient condition).
If $A\subseteq V$ and Assumption 3.1 holds, then
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=P_{A}H(\pi_{A}),$
with the convention that the right-hand side is 0 when $P_{A}=0$.
If $A\subseteq V$, then each $a\in A$ itself belongs to $R_{\text{Cn}}^{(V)}(a)$, so (H1) holds. Assumption 3.1 implies (H2) after restriction to $V$. Apply Theorem 6.5. ∎
Remark 6.7 (Meaning of the heterogeneous extension).
Theorem 6.5 shows that vocabulary heterogeneity matters only through the availability of zero-distortion representatives for core symbols. If every core element still has a distinct closure-equivalent representative in the receiver’s alphabet, then the exact zero-distortion rate is unchanged. If even one positive-probability core element has no such representative, zero distortion becomes impossible.
VII Conclusion
This paper studied lossy compression of a finite knowledge source equipped with a fixed proof system, with fidelity measured by preservation of deductive closure rather than by symbolwise equality. The central conclusion is that, under closure fidelity, the effective source complexity is governed by the irredundant core of the knowledge base rather than by the full stored source. Under an explicit disjointness condition on zero-distortion reconstruction sets, we showed that the zero-distortion rate is exactly $P_{A}H(\pi_{A})$, and that the full rate–distortion function decomposes into a contribution from that core alone whenever the reconstruction alphabet is contained in $\text{Cn}(S_{O})$. We further derived an exact rate–depth–distortion characterization under a bounded inference-depth budget; under an additional order-robustness assumption ensuring that the chosen core coincides with the essential set, this yields a sharp interpolation between classical symbolwise compression and unconstrained deductive compression.
From an information-theoretic perspective, the paper contributes a source-model formulation for deductive compression rather than merely a semantic communication architecture. Its main insight is that shared inference structure can be treated as a rate-reducing resource: redundant stored states are free under closure fidelity because they can be re-derived at the decoder. Source–channel separation statements, closure-adapted converse bounds, and heterogeneous-receiver extensions follow from this source-model viewpoint.
Several directions remain open. The most important is to remove or weaken the disjointness assumption and obtain a characterization for the general confusable case. Other natural directions are to extend the theory beyond finite-alphabet and i.i.d. models, to study multiterminal and interactive versions of deductive compression, and to allow heterogeneous proof systems across agents. We hope that the formulation developed here provides a useful starting point for a broader information-theoretic theory of structured sources with inference-enabled reconstruction.
Acknowledgment
The author thanks Associate Professor Rui Wang of the School of Computer Science at Shanghai Jiao Tong University for insightful comments, and doctoral students Yiming Wang, Chun Li, Hu Xu, Siyuan Qiu, Zeyan Li, Jiashuo Zhang, Junxuan He, and Xiao Wang for stimulating discussions throughout the writing and revision of this paper.
Appendix A Proofs of Sections III–VI
This appendix collects deferred proofs for the main results stated in Sections III–VI of the paper.
A-A Proof of Theorem 3.1
If $P_{A}=0$, then the claim follows from Remark 3.2, so assume henceforth that $P_{A}>0$.
The achievability part is Proposition 3.1, so it remains to prove the converse.
Let $P_{\hat{S}\mid S}$ be any conditional distribution achieving zero expected closure distortion. Define
$T(s_{o}):=\begin{cases}s_{o},&s_{o}\in A,\\ *,&s_{o}\in J.\end{cases}$
Since $T$ is a deterministic function of $S_{o}$, the data processing inequality gives
$I(S_{o};\hat{S}_{o})\geq I(T;\hat{S}_{o}).$
For each $a\in A$, the zero-distortion constraint implies that $P_{\hat{S}\mid S}(\cdot\mid a)$ is supported on $R_{\text{Cn}}(a)$. By Assumption 3.1, these supports are pairwise disjoint across $a\in A$. For $j\in J$, no analogous distinguishability constraint is required because redundant states are free under Proposition 2.2.
Write
$P_{a}:=P_{\hat{S}\mid S}(\cdot\mid a),\qquad\bar{\pi}:=\sum_{a\in A}\pi_{A}(a)\,P_{a}.$
Since the supports of $\{P_{a}\}_{a\in A}$ are pairwise disjoint, the disjoint-support entropy identity yields
$H(\bar{\pi})=H(\pi_{A})+\sum_{a\in A}\pi_{A}(a)\,H(P_{a}).$
If $P_{J}>0$, then define
$Q:=\sum_{j\in J}\frac{P_{O}(j)}{P_{J}}\,P_{\hat{S}\mid S}(\cdot\mid j).$
The output marginal is $P_{A}\bar{\pi}+P_{J}Q$, and concavity of entropy gives
$H(P_{A}\bar{\pi}+P_{J}Q)\geq P_{A}H(\bar{\pi})+P_{J}H(Q).$
Hence
$I(T;\hat{S}_{o})$ $=H(P_{A}\bar{\pi}+P_{J}Q)-P_{A}\sum_{a\in A}\pi_{A}(a)H(P_{a})-P_{J}H(Q)$
$\geq P_{A}\Bigl(H(\bar{\pi})-\sum_{a\in A}\pi_{A}(a)H(P_{a})\Bigr)$
$=P_{A}\,H(\pi_{A}).$
If $P_{J}=0$, then $P_{A}=1$ and the output marginal is simply $\bar{\pi}$. Therefore
$I(T;\hat{S}_{o})=H(\bar{\pi})-\sum_{a\in A}\pi_{A}(a)H(P_{a})=H(\pi_{A})=P_{A}H(\pi_{A}).$
Thus in all cases
$I(S_{o};\hat{S}_{o})\geq I(T;\hat{S}_{o})\geq P_{A}H(\pi_{A}).$
Combining this with Proposition 3.1 proves
$R_{\mathrm{sem}}(0;\,d_{\text{Cn}},\,P_{O})=P_{A}H(\pi_{A}).$
A-B Proof of Theorem 4.1
If $P_{A}=0$, then the source is supported on $J=S_{O}\setminus A$. Because $\hat{S}_{O}$ is nonempty and satisfies $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, choose any fixed $\hat{s}^{\star}\in\hat{S}_{O}$. For every $j\in J$, Proposition 2.2 gives
$d_{\text{Cn}}(j,\hat{s}^{\star}\mid S_{O})=0.$
Hence the constant reconstruction $\hat{S}_{o}\equiv\hat{s}^{\star}$ achieves zero distortion almost surely and therefore zero rate. Thus both sides of (4.2) equal 0. Assume $P_{A}>0$.
Step 1: distortion depends only on the core. Since $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, Proposition 2.2 implies
$d_{\text{Cn}}(j,\hat{s}_{o}\mid S_{O})=0,\qquad\forall\,j\in J,\ \forall\,\hat{s}_{o}\in\hat{S}_{O}.$
Therefore, for every test channel $P_{\hat{S}\mid S}$,
$\mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})]=P_{A}\,\mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})\mid S_{o}\in A].$
Thus the global distortion constraint $\mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})]\leq D$ is equivalent to
$\mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})\mid S_{o}\in A]\leq\frac{D}{P_{A}}.$
Step 2: lower bound on mutual information. Define
$T(s_{o}):=\begin{cases}s_{o},&s_{o}\in A,\\ *,&s_{o}\in J.\end{cases}$
Then $T$ is a deterministic function of $S_{o}$, so
$I(S_{o};\hat{S}_{o})\geq I(T;\hat{S}_{o}).$
Let $P_{a}:=P_{\hat{S}\mid S}(\cdot\mid a)$ for $a\in A$, and let
$\bar{P}_{A}:=\sum_{a\in A}\pi_{A}(a)\,P_{a}.$
The quantity
$H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})$
is exactly $I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o})$ for the core sub-source $\mathsf{A}_{o}\sim\pi_{A}$ and test channel $P_{\hat{S}\mid A}$.
If $0<P_{A}<1$, let
$Q:=\sum_{j\in J}\frac{P_{O}(j)}{P_{J}}\,P_{\hat{S}\mid S}(\cdot\mid j).$
Then the output marginal is
$P_{\hat{S}}=P_{A}\bar{P}_{A}+P_{J}Q.$
By concavity of entropy,
$H(P_{\hat{S}})\geq P_{A}H(\bar{P}_{A})+P_{J}H(Q).$
Also,
$H(\hat{S}_{o}\mid T)=P_{A}\sum_{a\in A}\pi_{A}(a)H(P_{a})+P_{J}H(Q).$
Hence
$I(T;\hat{S}_{o})$ $=H(P_{\hat{S}})-H(\hat{S}_{o}\mid T)$
$\geq P_{A}\Bigl(H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})\Bigr)$
$=P_{A}\,I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o}).$
If $P_{A}=1$, then $J=\varnothing$ and $T=S_{o}$ almost surely. The output marginal is simply $\bar{P}_{A}$, so
$I(T;\hat{S}_{o})$ $=H(\bar{P}_{A})-\sum_{a\in A}\pi_{A}(a)H(P_{a})$
$=I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o})=P_{A}\,I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o}).$
Therefore every feasible global test channel satisfies
$I(S_{o};\hat{S}_{o})\geq P_{A}\,R^{(A)}\!\Bigl(\frac{D}{P_{A}};\,d_{\text{Cn}},\,\pi_{A}\Bigr).$
Step 3: achievability by homogenizing the redundant part. Let $P_{\hat{S}\mid A}^{*}$ achieve $R^{(A)}(D/P_{A};\,d_{\text{Cn}},\,\pi_{A})$, and let
$\bar{P}_{A}^{*}:=\sum_{a\in A}\pi_{A}(a)\,P_{\hat{S}\mid A}^{*}(\cdot\mid a)$
be the corresponding output marginal. Extend $P_{\hat{S}\mid A}^{*}$ to a test channel on all of $S_{O}$ by
$P_{\hat{S}\mid S}(\cdot\mid s_{o}):=\begin{cases}P_{\hat{S}\mid A}^{*}(\cdot\mid s_{o}),&s_{o}\in A,\\[3.0pt] \bar{P}_{A}^{*}(\cdot),&s_{o}\in J.\end{cases}$
Because every $j\in J$ is distortion-free against every $\hat{s}_{o}\in\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, this extension preserves the distortion level $D$.
Moreover, the overall output marginal is exactly $\bar{P}_{A}^{*}$, so
$H(\hat{S}_{o})=H(\bar{P}_{A}^{*}),$
and
$H(\hat{S}_{o}\mid S_{o})=P_{A}H(\hat{S}_{o}\mid A_{o})+P_{J}H(\bar{P}_{A}^{*}).$
Therefore
$I(S_{o};\hat{S}_{o})$ $=H(\bar{P}_{A}^{*})-P_{A}H(\hat{S}_{o}\mid A_{o})-P_{J}H(\bar{P}_{A}^{*})$
$=P_{A}\Bigl(H(\bar{P}_{A}^{*})-H(\hat{S}_{o}\mid A_{o})\Bigr)$
$=P_{A}\,I(\mathsf{A}_{o};\hat{\mathsf{S}}_{o}),$
which equals
$P_{A}\,R^{(A)}\!\Bigl(\frac{D}{P_{A}};\,d_{\text{Cn}},\,\pi_{A}\Bigr).$
This proves (4.2). ∎
A-C Proof of Theorem 5.1
Fix $\delta\geq 0$.
If $P_{\delta}=0$, then the source is supported on $J_{\delta}:=S_{O}\setminus A_{\delta}$, and Lemma 5.1 implies that every $j\in J_{\delta}$ is zero-distortion free against every reconstruction symbol. Hence any constant reconstruction in $\hat{S}_{O}$ achieves zero $\delta$-distortion almost surely, so
$R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})=0,$
which agrees with the convention $P_{\delta}H(\pi_{\delta})=0$.
Assume henceforth that $P_{\delta}>0$, and write $J_{\delta}:=S_{O}\setminus A_{\delta}$. Let $P_{\hat{S}\mid S}$ be any test channel achieving zero expected $\delta$-distortion. Define
$T_{\delta}(s_{o}):=\begin{cases}s_{o},&s_{o}\in A_{\delta},\\ *,&s_{o}\in J_{\delta}.\end{cases}$
Since $T_{\delta}$ is a deterministic function of $S_{o}$,
$I(S_{o};\hat{S}_{o})\geq I(T_{\delta};\hat{S}_{o}).$
For each $a\in A_{\delta}$, zero $\delta$-distortion implies that $P_{\hat{S}\mid S}(\cdot\mid a)$ is supported on $R_{\delta}(a)$. By Assumption 5.3, these supports are pairwise disjoint across $a\in A_{\delta}$. For $j\in J_{\delta}$, the output is unconstrained by Lemma 5.1.
Write
$P_{a}:=P_{\hat{S}\mid S}(\cdot\mid a),\qquad\bar{\pi}_{\delta}:=\sum_{a\in A_{\delta}}\pi_{\delta}(a)\,P_{a}.$
Since the supports of $\{P_{a}\}_{a\in A_{\delta}}$ are pairwise disjoint, the disjoint-support entropy identity gives
$H(\bar{\pi}_{\delta})=H(\pi_{\delta})+\sum_{a\in A_{\delta}}\pi_{\delta}(a)H(P_{a}).$
If $0<P_{\delta}<1$, define
$Q_{\delta}:=\sum_{j\in J_{\delta}}\frac{P_{O}(j)}{1-P_{\delta}}\,P_{\hat{S}\mid S}(\cdot\mid j).$
Then the output marginal is
$P_{\hat{S}}=P_{\delta}\bar{\pi}_{\delta}+(1-P_{\delta})Q_{\delta}.$
By concavity of entropy,
$H(P_{\hat{S}})\geq P_{\delta}H(\bar{\pi}_{\delta})+(1-P_{\delta})H(Q_{\delta}).$
Also,
$H(\hat{S}_{o}\mid T_{\delta})=P_{\delta}\sum_{a\in A_{\delta}}\pi_{\delta}(a)H(P_{a})+(1-P_{\delta})H(Q_{\delta}).$
Hence
$I(T_{\delta};\hat{S}_{o})$ $=H(P_{\hat{S}})-H(\hat{S}_{o}\mid T_{\delta})$
$\geq P_{\delta}\Bigl(H(\bar{\pi}_{\delta})-\sum_{a\in A_{\delta}}\pi_{\delta}(a)H(P_{a})\Bigr)$
$=P_{\delta}H(\pi_{\delta}).$
If $P_{\delta}=1$, then $J_{\delta}=\varnothing$ and $T_{\delta}=S_{o}$ almost surely. The output marginal is simply $\bar{\pi}_{\delta}$, so
$I(T_{\delta};\hat{S}_{o})$ $=H(\bar{\pi}_{\delta})-\sum_{a\in A_{\delta}}\pi_{\delta}(a)H(P_{a})$
$=H(\pi_{\delta})=P_{\delta}H(\pi_{\delta}).$
Therefore, in all cases with $P_{\delta}>0$,
$I(S_{o};\hat{S}_{o})\geq I(T_{\delta};\hat{S}_{o})\geq P_{\delta}H(\pi_{\delta}).$
For achievability, define
$P_{\hat{S}\mid S}(\hat{s}_{o}\mid s_{o}):=\begin{cases}\mathbf{1}[\hat{s}_{o}=s_{o}],&s_{o}\in A_{\delta},\\[3.0pt] \pi_{\delta}(\hat{s}_{o}),&s_{o}\in J_{\delta},\end{cases}$
with $\pi_{\delta}$ supported on $A_{\delta}\subseteq\hat{S}_{O}$. If $s_{o}\in A_{\delta}$, the reconstruction is exact and hence has zero $\delta$-distortion. If $s_{o}\in J_{\delta}$, zero $\delta$-distortion follows from Lemma 5.1. The output marginal is $\pi_{\delta}$, and exactly as in the proof of Proposition 3.1,
$I(S_{o};\hat{S}_{o})=P_{\delta}H(\pi_{\delta}).$
Thus
$R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})=P_{\delta}H(\pi_{\delta}).$
A-D Proof of Corollary 5.1
Part (i) follows from Theorem 5.1 at $\delta=0$. Indeed, $A_{0}=S_{O}$, so $P_{0}=1$ and $\pi_{0}=P_{O}$, which gives
$R_{\mathrm{sem}}(0,0;\,d_{\text{Cn}}^{0},\,P_{O})=H(P_{O}).$
Part (ii) follows from Proposition 5.1(iii) and Theorem 5.1: for every $\delta\geq\mathsf{D_{d}}$, one has $A_{\delta}=A$, hence $P_{\delta}=P_{A}$ and $\pi_{\delta}=\pi_{A}$.
For part (iii), note that if $\delta_{1}\leq\delta_{2}$, then
$d_{\text{Cn}}^{\delta_{2}}(s_{o},\hat{s}_{o}\mid S_{O})\leq d_{\text{Cn}}^{\delta_{1}}(s_{o},\hat{s}_{o}\mid S_{O})$
for all $(s_{o},\hat{s}_{o})$, because allowing more iterations (of $T_{\mathsf{PS}}$) can only enlarge the zero-distortion feasible set. Therefore the feasible test-channel set for zero $\delta_{2}$-distortion contains that for zero $\delta_{1}$-distortion, and thus
$R_{\mathrm{sem}}(0,\delta_{2};\,d_{\text{Cn}}^{\delta_{2}},\,P_{O})\leq R_{\mathrm{sem}}(0,\delta_{1};\,d_{\text{Cn}}^{\delta_{1}},\,P_{O}).$
So $\delta\mapsto R_{\mathrm{sem}}(0,\delta)$ is non-increasing. ∎
A-E Proof of Theorem 5.2
Fix $\delta\geq 0$. If $P_{\delta}=0$, then the source is supported on $S_{O}\setminus A_{\delta}$. Since $\hat{S}_{O}$ is nonempty, any constant reconstruction in $\hat{S}_{O}$ has zero $\delta$-distortion almost surely by Lemma 5.1. Hence
$R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})=0,\qquad\forall\,D\geq 0,$
which agrees with the convention in (5.5). Assume henceforth that $P_{\delta}>0$.
By Lemma 5.1, every $j\in S_{O}\setminus A_{\delta}$ is distortion-free against every reconstruction symbol under $d_{\text{Cn}}^{\delta}$. Therefore, for every test channel,
$\mathbb{E}[d_{\text{Cn}}^{\delta}(S_{o},\hat{S}_{o})]=P_{\delta}\,\mathbb{E}[d_{\text{Cn}}^{\delta}(S_{o},\hat{S}_{o})\mid S_{o}\in A_{\delta}].$
Thus the distortion constraint $\mathbb{E}[d_{\text{Cn}}^{\delta}]\leq D$ reduces to a core-only constraint with threshold $D/P_{\delta}$.
Now define
$T_{\delta}(s_{o}):=\begin{cases}s_{o},&s_{o}\in A_{\delta},\\ *,&s_{o}\notin A_{\delta}.\end{cases}$
As in the proof of Theorem 4.1,
$I(S_{o};\hat{S}_{o})\geq I(T_{\delta};\hat{S}_{o})\geq P_{\delta}\,I(\mathsf{A}_{\delta,o};\hat{\mathsf{S}}_{o}),$
where $\mathsf{A}_{\delta,o}\sim\pi_{\delta}$ and the channel is the restriction of $P_{\hat{S}\mid S}$ to $A_{\delta}$. Taking the minimum over all feasible channels yields
$R_{\mathrm{sem}}(D,\delta)\geq P_{\delta}\,R^{(A_{\delta})}\!\Bigl(\frac{D}{P_{\delta}};\,d_{\text{Cn}}^{\delta},\,\pi_{\delta}\Bigr).$
For achievability, let $P_{\hat{S}\mid A_{\delta}}^{*}$ achieve the right-hand side, and extend it to all of $S_{O}$ by assigning to every $j\notin A_{\delta}$ the output marginal induced by $P_{\hat{S}\mid A_{\delta}}^{*}$. This keeps the distortion unchanged because the symbols outside $A_{\delta}$ are $\delta$-redundant, and the same entropy calculation as in Appendix A-B shows that the resulting mutual information is exactly
$P_{\delta}\,R^{(A_{\delta})}\!\Bigl(\frac{D}{P_{\delta}};\,d_{\text{Cn}}^{\delta},\,\pi_{\delta}\Bigr).$
Hence (5.5) holds. ∎
A-F Proof of Theorem 6.1
All distortion functions used in the theorem are bounded single-letter distortions on finite alphabets. Therefore the standard source–channel separation theorem applies.
For closure distortion $d_{\text{Cn}}$, the source-coding side is governed by $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$, and the channel-coding side is governed by the Shannon capacity $C(W)$. Hence a distortion level $D$ is achievable whenever
$R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})<\kappa C(W),$
and conversely every achievable distortion level must satisfy
$R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})\leq\kappa C(W).$
The same reasoning applies to the bounded-inference distortion $d_{\text{Cn}}^{\delta}$: for every fixed $\delta$,
$R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})<\kappa C(W)$
is sufficient for achievability, whereas
$R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})\leq\kappa C(W)$
is necessary.
Finally, under the hypotheses of Theorems 3.1 and 5.1, substituting the exact zero-distortion formulas gives the displayed sufficient and necessary conditions (6.3)–(6.6). ∎
A-G Proof of Theorem 6.4
If $A=\varnothing$, then by Proposition 2.2 every input is closure-correct against every $\hat{s}_{o}\in\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, so $\epsilon_{\text{Cn}}=0$ and the claim is trivial. Assume henceforth that $A\neq\varnothing$.
Because $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, Proposition 2.2 implies that
$d_{\text{Cn}}(j,\hat{s}_{o}\mid S_{O})=0,\qquad\forall\,j\in J,\ \forall\,\hat{s}_{o}\in\hat{S}_{O}.$
Hence closure errors can arise only on core inputs, and therefore
$\epsilon_{\text{Cn}}=\sum_{a\in A}P_{O}(a)\,\Pr[\hat{S}_{o}\notin R_{\text{Cn}}(a)\mid S_{o}=a].$
Let $B:=\mathbf{1}[S_{o}\in A]$. Since $B$ is a deterministic function of $S_{o}$,
$I(S_{o};\hat{S}_{o})=I(B,S_{o};\hat{S}_{o})=I(B;\hat{S}_{o})+I(S_{o};\hat{S}_{o}\mid B)\geq I(S_{o};\hat{S}_{o}\mid B).$
Therefore
$I(S_{o};\hat{S}_{o})\geq P_{A}I(S_{o};\hat{S}_{o}\mid B=1),$
since the conditional mutual information given $B=0$ is nonnegative.
Conditioned on $B=1$, the source is distributed as $\pi_{A}$ on the alphabet $A$. If $|A|=1$, then $H(\pi_{A})=0$, so the desired inequality reduces to the trivial bound $I(S_{o};\hat{S}_{o})\geq 0$. Assume now that $|A|\geq 2$.
Under Assumption 3.1, define $g:\hat{S}_{O}\to A\cup\{\mathsf{e}\}$ by
[…]
Because the sets $R_{\text{Cn}}(a)$ are pairwise disjoint, $g$ is well-defined. By data processing,
$H(S_{o}\mid\hat{S}_{o},\ B=1)\leq H(S_{o}\mid g(\hat{S}_{o}),\ B=1).$
Let $p_{A}:=\epsilon_{\text{Cn}}/P_{A}$. Applying the classical Fano inequality [7, Sec. 2.10][22, Sec. 1.9] to the $|A|$-ary source conditioned on $B=1$ gives
$H(S_{o}\mid g(\hat{S}_{o}),\ B=1)\leq h_{b}(p_{A})+p_{A}\log(|A|-1).$
Hence
$I(S_{o};\hat{S}_{o}\mid B=1)\geq H(\pi_{A})-h_{b}(p_{A})-p_{A}\log(|A|-1).$
Multiplying by $P_{A}$ yields
$I(S_{o};\hat{S}_{o})\geq P_{A}H(\pi_{A})-P_{A}h_{b}\!\Bigl(\frac{\epsilon_{\text{Cn}}}{P_{A}}\Bigr)-\epsilon_{\text{Cn}}\log(|A|-1).$
Finally, Jensen’s inequality gives
$P_{A}h_{b}\!\Bigl(\frac{\epsilon_{\text{Cn}}}{P_{A}}\Bigr)\leq h_{b}(\epsilon_{\text{Cn}}),$
which proves
$I(S_{o};\hat{S}_{o})\geq P_{A}H(\pi_{A})-h_{b}(\epsilon_{\text{Cn}})-\epsilon_{\text{Cn}}\log(|A|-1).$
A-H Proof of Theorem 6.5
If $P_{A}=0$, then the source is supported on $J=S_{O}\setminus A$. Since $A\neq\varnothing$ and (H1) holds, choose any $a_{0}\in A$ and any representative $\phi(a_{0})\in R_{\text{Cn}}^{(V)}(a_{0})$. Because
$d_{\text{Cn}}(a_{0},\phi(a_{0})\mid S_{O})=0,$
one has
$\text{Cn}((S_{O}\setminus\{a_{0}\})\cup\{\phi(a_{0})\})=\text{Cn}(S_{O}),$
and hence by reflexivity $\phi(a_{0})\in\text{Cn}(S_{O})$. Therefore Proposition 2.2 gives
$d_{\text{Cn}}(j,\phi(a_{0})\mid S_{O})=0,\qquad\forall\,j\in J.$
So the constant reconstruction $\hat{S}_{o}\equiv\phi(a_{0})$ achieves zero distortion almost surely, and hence
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=0,$
which agrees with the stated convention. Assume henceforth that $P_{A}>0$.
For each $a\in A$, assumption (H1) ensures that $R_{\text{Cn}}^{(V)}(a)\neq\varnothing$, so choose one representative $\phi(a)\in R_{\text{Cn}}^{(V)}(a)$. By (H2), the map $\phi:A\to V$ is injective.
Define a test channel by
$P_{\hat{S}\mid S}(\hat{s}\mid s):=\begin{cases}\mathbf{1}[\hat{s}=\phi(s)],&s\in A,\\[3.0pt] \displaystyle\sum_{a\in A}\pi_{A}(a)\,\mathbf{1}[\hat{s}=\phi(a)],&s\in J.\end{cases}$
For $a\in A$, zero distortion holds because $\phi(a)\in R_{\text{Cn}}^{(V)}(a)$ by construction. For $j\in J$, every possible output is of the form $\phi(a)$ with $a\in A$. Since
$d_{\text{Cn}}(a,\phi(a)\mid S_{O})=0,$
the equality of closures implies
$\text{Cn}((S_{O}\setminus\{a\})\cup\{\phi(a)\})=\text{Cn}(S_{O}),$
and hence by reflexivity $\phi(a)\in\text{Cn}(S_{O})$. Therefore Proposition 2.2 gives
$d_{\text{Cn}}(j,\phi(a)\mid S_{O})=0.$
So the expected distortion is zero.
Because $\phi$ is injective, the induced output distribution on $\phi(A)$ has entropy $H(\pi_{A})$, and exactly the same calculation as in Proposition 3.1 yields
$I(S_{o};\hat{S}_{o})=P_{A}H(\pi_{A}).$
Hence
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})\leq P_{A}H(\pi_{A}).$
For the converse, let $P_{\hat{S}\mid S}$ be any feasible zero-distortion test channel supported on $V$. Then for every $a\in A$, the conditional law $P_{\hat{S}\mid S}(\cdot\mid a)$ must be supported on $R_{\text{Cn}}^{(V)}(a)$. By assumption (H2), these supports are pairwise disjoint, so the converse proof of Theorem 3.1 applies verbatim with $R_{\text{Cn}}(a)$ replaced by $R_{\text{Cn}}^{(V)}(a)$. Thus
$I(S_{o};\hat{S}_{o})\geq P_{A}H(\pi_{A}),$
and therefore
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=P_{A}H(\pi_{A}).$
For the final claim, if some $a\in A$ with $P_{O}(a)>0$ satisfies $R_{\text{Cn}}^{(V)}(a)=\varnothing$, then there is no reconstruction symbol in $V$ that can reproduce $a$ with zero closure distortion. Hence the zero-distortion feasible set is empty, and
$R_{\mathrm{sem}}^{(V)}(0;\,d_{\text{Cn}},\,P_{O})=\infty.$
Appendix B Datalog Classes Satisfying the Assumptions
This appendix records concrete Datalog classes under which the standing assumptions used in the paper are satisfied. The goal is not to give the most general logical characterization, but to isolate natural and checkable sufficient conditions.
B-A Finite Active-Domain Datalog
Definition B.1 (Finite active-domain Datalog instance).
A finite active-domain Datalog instance consists of:
  1. (i)
    a function-free Datalog program $P$;
  2. (ii)
    a finite active domain $\mathcal{D}$;
  3. (iii)
    the finite set $\mathbb{S}_{O}$ of all ground atoms over $\mathcal{D}$ whose predicates occur in $P$;
  4. (iv)
    a finite stored knowledge base $S_{O}\subseteq\mathbb{S}_{O}$.
The proof system is the standard Datalog immediate-consequence semantics induced by $P$.
Remark B.1.
These properties are standard consequences of least-fixpoint Datalog semantics over finite active domains [29, 5, 1, 9].
Proposition B.1 (Basic assumptions hold for finite active-domain Datalog).
For every finite active-domain Datalog instance, Assumptions 2.12.3 and Assumption 5.1 hold.
Because the program is function-free and the active domain is finite, the set $\mathbb{S}_{O}$ of ground atoms is finite. Hence every knowledge base $S_{O}\subseteq\mathbb{S}_{O}$ is finite and effectively listable.
Derivability in Datalog is decidable by iterating the immediate consequence operator $T_{P}$ until the least fixpoint is reached [29, 5, 1, 9]. Since the active domain is finite, the chain of iterates stabilizes in finitely many steps. Therefore membership $s\in\text{Cn}(\Gamma)$ is decidable for every finite $\Gamma\subseteq\mathbb{S}_{O}$ and every $s\in\mathbb{S}_{O}$, proving Assumptions 2.1 and 2.3.
Finally, for Assumption 5.1, one may take the inflationary immediate-consequence operator
$T_{\mathsf{PS}}(B):=B\cup T_{P}(B),$
where $T_{P}$ denotes the standard one-step Datalog consequence operator induced by $P$. This operator is monotone, inflationary, computable, and over a finite active domain its iterates stabilize in finitely many steps at the least fixpoint, which equals the deductive closure. Thus Assumption 5.1 holds as well. ∎
B-B A Concrete Class Ensuring Core-Disjointness
Definition B.2 (Tag-seeded Datalog class).
Let $A=\text{Atom}(S_{O})$. A finite active-domain Datalog instance is called tag-seeded if there exists an injective map $\tau:A\to\mathcal{D}$ assigning a unique tag to each core fact, together with a distinguished argument position in the relevant predicates, such that:
  1. (TS1)
    each $a\in A$ appears with tag $\tau(a)$, and no other stored fact in $S_{O}$ carries the tag $\tau(a)$;
  2. (TS2)
    every instantiated rule propagates but never creates the distinguished tag: if all body facts used in a derivation carry the same tag $t$, then the derived head fact also carries tag $t$; moreover, a fact carrying tag $t$ can be derived only from premises of which at least one already carries tag $t$;
  3. (TS3)
    for every $a\in A$, there exists at least one fact $w_{a}\in\text{Cn}(S_{O})$ carrying tag $\tau(a)$ such that $w_{a}\notin\text{Cn}(S_{O}\setminus\{a\})$.
Proposition B.2 (Tag seeding implies core-disjointness).
Let a finite active-domain Datalog instance be tag-seeded in the sense of Definition B.2, and let the reconstruction alphabet satisfy $\hat{S}_{O}\subseteq\text{Cn}(S_{O})\cap\mathbb{S}_{O}$. Then Assumption 3.1 holds.
Fix $a\in A$, and let $\hat{s}\in R_{\text{Cn}}(a)$. Then
$\text{Cn}\bigl((S_{O}\setminus\{a\})\cup\{\hat{s}\}\bigr)=\text{Cn}(S_{O}).$
In particular, the witness fact $w_{a}$ belongs to the left-hand closure. By (TS2), any derivation of $w_{a}$ must preserve the tag $\tau(a)$. By (TS1), the knowledge base $S_{O}\setminus\{a\}$ contains no stored fact carrying the tag $\tau(a)$. Therefore $\hat{s}$ itself must carry the tag $\tau(a)$.
Hence every element of $R_{\text{Cn}}(a)$ carries the tag $\tau(a)$. Because the tags are distinct across different core facts by injectivity of $\tau$, no reconstruction symbol can belong simultaneously to $R_{\text{Cn}}(a_{1})$ and $R_{\text{Cn}}(a_{2})$ when $a_{1}\neq a_{2}$. Thus
$R_{\text{Cn}}(a_{1})\cap R_{\text{Cn}}(a_{2})=\varnothing,\qquad a_{1}\neq a_{2},$
which is Assumption 3.1. ∎
Remark B.2 (Finite-depth variants).
The tagged-provenance construction above justifies Assumption 3.1 for the unconstrained zero-distortion problem. A finite-depth analogue can also be imposed to justify Assumption 5.3, but it requires tagging the $\delta$-irredundant symbols themselves rather than only the final core $A$. Since this extra bookkeeping is orthogonal to the main line of the paper, we do not formalize the finite-depth syntactic condition here.
Remark B.3 (Scope of the syntactic condition).
The tag-seeded class is only a sufficient condition for Assumptions 3.1 and 5.3, not a necessary one. Its role is to show that the disjointness assumptions used in the main theorems hold on a concrete and checkable Datalog family, rather than being purely abstract hypotheses.
Appendix C Extra Examples and Counterexamples
This appendix records supplementary examples illustrating the main definitions and the limits of the theory.
C-A Order Dependence of the Irredundant Core
Example C.1 (Order dependence).
Let
$S_{O}=\{p,q,r\},$
and suppose the proof system contains the two rules
$p\leftarrow q,\qquad q\leftarrow p.$
Then $p$ and $q$ are deductively interchangeable, while $r$ is independent.
If the canonical order is $p\prec q\prec r$, then $p$ is deleted when scanned because $p\in\text{Cn}(\{q,r\})$, and the resulting core is
$\text{Atom}(S_{O})=\{q,r\}.$
If the canonical order is $q\prec p\prec r$, then the resulting core is
$\text{Atom}(S_{O})=\{p,r\}.$
Thus $\text{Atom}(S_{O})$ may depend on the canonical order, even though in both cases the closure it generates is the same.
C-B A sufficient condition for order invariance
Definition C.1 (Essential set).
The essential set of $S_{O}$ is
$\operatorname{Ess}(S_{O}):=\{s\in S_{O}:\ s\notin\text{Cn}(S_{O}\setminus\{s\})\}.$
Proposition C.1 (Order-robust core under essential generation).
For every finite knowledge base $S_{O}$:
  1. (i)
    $\operatorname{Ess}(S_{O})$ depends only on $S_{O}$ and Cn, not on the canonical order;
  2. (ii)
    $\operatorname{Ess}(S_{O})\subseteq\text{Atom}(S_{O})$ for every canonical order;
  3. (iii)
    if $\text{Cn}(\operatorname{Ess}(S_{O}))=\text{Cn}(S_{O})$, then
    $\text{Atom}(S_{O})=\operatorname{Ess}(S_{O})$
    for every canonical order.
Part (i) is immediate from the definition.
For (ii), let $s\in\operatorname{Ess}(S_{O})$, and let $B$ be the current set when $s$ is scanned in the deletion procedure of Definition 2.2. Since $B\setminus\{s\}\subseteq S_{O}\setminus\{s\}$, monotonicity of Cn gives
$\text{Cn}(B\setminus\{s\})\subseteq\text{Cn}(S_{O}\setminus\{s\}).$
Because $s\notin\text{Cn}(S_{O}\setminus\{s\})$, one also has $s\notin\text{Cn}(B\setminus\{s\})$, so $s$ cannot be deleted. Hence $s\in\text{Atom}(S_{O})$.
For (iii), assume $\text{Cn}(\operatorname{Ess}(S_{O}))=\text{Cn}(S_{O})$ and let $s\in\text{Atom}(S_{O})$. If $s\notin\operatorname{Ess}(S_{O})$, then $s\in\text{Cn}(S_{O}\setminus\{s\})$, while by part (ii) all essential elements survive every deletion order. Hence, at the moment $s$ is scanned, the current set still contains $\operatorname{Ess}(S_{O})\setminus\{s\}=\operatorname{Ess}(S_{O})$. By monotonicity,
$\text{Cn}(\operatorname{Ess}(S_{O}))\subseteq\text{Cn}(B\setminus\{s\}).$
Since $\text{Cn}(\operatorname{Ess}(S_{O}))=\text{Cn}(S_{O})$ by assumption and $s\in\text{Cn}(S_{O})$ by reflexivity, it follows that
$s\in\text{Cn}(B\setminus\{s\}),$
which would force $s$ to be deleted—a contradiction. Therefore every element of $\text{Atom}(S_{O})$ is essential, so together with part (ii) one gets $\text{Atom}(S_{O})=\operatorname{Ess}(S_{O})$. ∎
Corollary C.1 (Order-invariant materialized EDB/IDB stores).
Suppose $S_{O}$ consists of EDB facts together with materialized IDB facts of a function-free Datalog program whose rule heads are all IDB predicates. Then no EDB fact is derivable from the remaining stored facts, the EDB part generates $\text{Cn}(S_{O})$, and therefore $\text{Atom}(S_{O})$ equals the stored EDB part for every canonical order.
Since all rule heads are IDB predicates, no rule derives an EDB fact. Hence every stored EDB fact is essential. Moreover, the stored EDB facts generate all materialized IDB facts by construction, so the EDB subset generates $\text{Cn}(S_{O})$. Apply Proposition C.1(iii). ∎
C-C A Minimal Exact Zero-Distortion Example
Example C.2 (One redundant consequence).
Let
$S_{O}=\{a,b,c\},$
and suppose $c\in\text{Cn}(\{a,b\})$, while neither $a$ nor $b$ is derivable from the other two statements. Then
$A=\text{Atom}(S_{O})=\{a,b\},\qquad J=\{c\}.$
Assume further that $S_{O}\subseteq\hat{S}_{O}$ and that Assumption 3.1 holds for this instance.
If $P_{O}$ is uniform on $S_{O}$, then
$P_{A}=\frac{2}{3},\qquad\pi_{A}(a)=\pi_{A}(b)=\frac{1}{2}.$
Hence Theorem 3.1 gives
$R_{\mathrm{sem}}(0)=P_{A}H(\pi_{A})=\frac{2}{3}\cdot 1=\frac{2}{3}\ \text{bits}.$
By contrast, classical zero-distortion coding under Hamming fidelity requires
$H(P_{O})=\log 3$
bits. Thus the redundant consequence $c$ creates a strict deductive compression gain.
C-D A Simple Rate–Depth Example
Example C.3 (A two-step shortcut).
Let
$S_{O}=\{a,b,d,e\},$
and suppose the proof system contains auxiliary statements $c,f\notin S_{O}$ and rules
$c\leftarrow a,\qquad d\leftarrow c,\qquad f\leftarrow b,\qquad e\leftarrow f.$
Then $d$ is derivable from $a$ with derivation depth $2$, and $e$ is derivable from $b$ with derivation depth $2$. The irredundant core is
$A=\{a,b\}.$
Assume throughout this example that $S_{O}\subseteq\hat{S}_{O}$, that Assumption 5.2 holds, and that Assumption 5.3 holds for $\delta=0,1,2$.
However, with depth budget $\delta=1$, neither $d$ nor $e$ is redundant relative to $S_{O}\setminus\{d\}$ or $S_{O}\setminus\{e\}$, because the intermediate statements $c$ and $f$ are not stored. Hence
$A_{0}=A_{1}=S_{O}.$
At depth budget $\delta=2$, both $d$ and $e$ become redundant, so
$A_{2}=\{a,b\}=A.$
If $P_{O}$ is uniform on $S_{O}$, then
$R_{\mathrm{sem}}(0,0)=R_{\mathrm{sem}}(0,1)=H(P_{O})=2$
bits, whereas
$R_{\mathrm{sem}}(0,2)=P_{A}H(\pi_{A})=\frac{1}{2}\cdot 1=\frac{1}{2}$
bit. This illustrates how additional inference depth can sharply reduce the effective source rate.
C-E A Confusable-Core Counterexample
Example C.4 (Failure of Assumption 3.1).
Let
$S_{O}=\{a_{1},a_{2},b\},$
and let the proof system contain an auxiliary statement $r\notin S_{O}$ together with rules
$r\leftarrow a_{1},a_{2},\qquad a_{1}\leftarrow b,r,\qquad a_{2}\leftarrow b,r.$
Neither $a_{1}$ nor $a_{2}$ is derivable from the remaining stored statements alone, so both are core elements.
Now replace $a_{1}$ by $r$. From $\{a_{2},b,r\}$, one can derive $a_{1}$, and the resulting closure equals $\text{Cn}(S_{O})$. Similarly, replacing $a_{2}$ by $r$ also preserves the full closure. Therefore
$r\in R_{\text{Cn}}(a_{1})\cap R_{\text{Cn}}(a_{2}),$
so Assumption 3.1 fails.
This example shows that the general zero-distortion problem has a real core-confusability phenomenon and is not captured by the disjoint case alone.
C-F Restricted Alphabet Impossibility
Example C.5 (No representative for a core symbol).
Let
$S_{O}=\{a,b\},$
and suppose the proof system contains no nontrivial derivation rules. Then
$A=\text{Atom}(S_{O})=\{a,b\}.$
If the reconstruction alphabet is restricted to
$V=\{a\},$
then $a$ has a zero-distortion representative in $V$, but $b$ does not:
$R_{\text{Cn}}^{(V)}(b)=\varnothing.$
Therefore zero closure distortion is impossible, and Theorem 6.5 yields
$R_{\mathrm{sem}}^{(V)}(0)=\infty.$
This is the simplest instance showing that heterogeneous receivers matter only through the availability of zero-distortion representatives for core symbols.
C-G Numerical Illustration on Materialized Datalog Stores
We briefly illustrate the zero-distortion law on a family of function-free Datalog stores consisting of EDB facts together with a materialized subset of derived IDB facts, a viewpoint closely related to the database literature on materialization and materialized views [12, 11, 6]. By Corollary C.1, the irredundant core is exactly the stored EDB part for every canonical order, so the example also illustrates an order-invariant regime of the theory.
For concreteness, the synthetic instance uses three EDB predicates $\mathsf{connected}$, $\mathsf{supplies}$, and $\mathsf{produces}$, and two IDB predicates $\mathsf{reachable}$ and $\mathsf{available}$, with rules
$\mathsf{reachable}(X,Y)\leftarrow\mathsf{connected}(X,Y),\qquad\mathsf{reachable}(X,Z)\leftarrow\mathsf{reachable}(X,Y),\,\mathsf{connected}(Y,Z),$
$\mathsf{available}(I,L)\leftarrow\mathsf{produces}(S,I),\,\mathsf{supplies}(S,L),$
$\mathsf{available}(I,L)\leftarrow\mathsf{produces}(S,I),\,\mathsf{supplies}(S,L_{0}),\,\mathsf{reachable}(L_{0},L).$
The numerical values below are illustrative only and play no role in any theorem. For the zero-distortion formula quoted below, assume also that $S_{O}\subseteq\hat{S}_{O}$ and that Assumption 3.1 holds on this synthetic family.
Consider a fixed 200-location supply-chain Datalog instance with $|A|=1705$ stored EDB facts. Let $S_{O}:=A\cup J_{\mu}$, where $J_{\mu}$ is a materialized subset of derived IDB facts of size controlled by a materialization level $\mu$. Under the uniform source on $S_{O}$, Theorem 3.1 gives
$R_{\mathrm{sem}}(0)=\frac{|A|}{|S_{O}|}\log|A|,\qquad R(0;d_{H})=\log|S_{O}|.$
Table I reports the resulting compression ratio $R_{\mathrm{sem}}(0)/R(0;d_{H})$.
TABLE I: Synthetic illustration of deductive compression gain under materialized redundant consequences (uniform source).
$\mu$ (%) $|J|$ $|S_{O}|$ $R_{\mathrm{sem}}(0)/R(0;d_{H})$ $\log|A|/\log|S_{O}|$
0 0 1 705 1.000 1.000
10 4 340 6 045 0.241 0.855
20 8 680 10 385 0.132 0.805
30 13 020 14 725 0.090 0.775
50 21 700 23 405 0.054 0.740
80 34 720 36 425 0.033 0.709
100 43 400 45 105 0.026 0.694
The table illustrates the basic mechanism of the theory: as more deductively redundant consequences are materialized and stored, the irredundant core remains unchanged while the classical zero-distortion benchmark grows with $|S_{O}|$. Accordingly, the closure-based zero-distortion rate becomes a rapidly shrinking fraction of the classical Hamming zero-distortion rate.